### v1

# Algebraic groups over free and hyperbolic groups

Abstract
We define an algebraic group over a group $G$ to be a variety - that is, a subset of $G^{d}$ defined by equations over $G$ endowed with a group law whose coordinates can be expressed as word maps. In the case where $G$ is a torsion-free hyperbolic group and the underlying variety is irreducible we give a complete description of all algebraic groups.
1 Introduction
In classical algebraic geometry, an algebraic group over a field $k$ is an algebraic variety endowed with a group law which can be expressed by polynomials in the coordinates. We ask what happens when we replace the field $k$ by a group $\Gamma$, and polynomials by word maps with constants in $\Gamma$.
More precisely, given a group $\Gamma$, an equation over $\Gamma$ in the variables $x_{1},\ldots,x_{d}$ is an expression of the form $w(x_{1},\ldots,x_{d})=1$, where $w$ is a word in the variables $x_{1}^{\pm},\dots,x_{d}^{\pm}$ with constants in $\Gamma$, i. e. an element of the free product $\Gamma*\langle x_{1},\ldots,x_{d}\rangle$. Given a set $\Sigma$ of such words, the variety $V$ defined by $\Sigma$ is the subset of $\Gamma^{d}$ consisting of the $d$-tuples $(v_{1},\dots,v_{d})\in\Gamma^{d}$ satisfying all the equations $w(v_{1},\dots,v_{d})=1$ for $w\in\Sigma$. An algebraic map $F:V\rightarrow V^{\prime}$ between two algebraic varieties $V\subset\Gamma^{d}$ and $V^{\prime}\subset\Gamma^{d^{\prime}}$ is a map such that there exist words $w_{1},\dots,w_{d^{\prime}}\in\Gamma*\langle x_{1},\dots,x_{d}\rangle$ such that for all $v=(v_{1},\dots,v_{d})\in V$, $F(v)=(w_{1}(v),\dots,w_{d^{\prime}}(v))$. An isomorphism of algebraic varieties $F:V\rightarrow V^{\prime}$ is a bijection such that $F$ and $F^{-1}$ are algebraic maps. When $\Gamma$ is a non-elementary torsion-free hyperbolic group, the set of algebraic varieties is the collection of closed sets for a topology, called the Zariski topology (the hypothesis on $\Gamma$ ensures that this is stable under finite unions, see [BMR99, §3.1]).
We define an algebraic group over $\Gamma$ as a group $V$ which is a variety and whose multiplication law $\mu:V\times V\to V$ is an algebraic map. It would be natural to ask additionally that the map $V\rightarrow V$ sending and element to its inverse is also an algebraic map, but we actually don’t require this as our results hold without this assumption (and in fact imply that this extra condition automatically holds).
Example 1.1.
Let $\Gamma$ be any group. Then $\Gamma$ itself is an algebraic group for its standard multiplication law $\mu(u,v)=uv$, which is simply the algebraic map $\Gamma^{2}\rightarrow\Gamma$ associated to the word $xy\in\Gamma*\langle x,y\rangle$. We note that the inverse $v\mapsto v^{-1}$ is also an algebraic map.
Example 1.2.
If $a\in\Gamma$ is any element, then its centralizer $V=Z_{\Gamma}(a)$ is an algebraic group with underlying variety defined by the equation $xax^{-1}a^{-1}=1$, and endowed with the standard multiplication law. Similarly, the centralizer of any subset of $\Gamma$ is an algebraic group. Now if $\Gamma$ is a torsion-free hyperbolic group, any centralizer is either the trivial group, the whole group $\Gamma$, or a maximal cyclic subgroup $\langle c\rangle=Z(c)$.
Example 1.3.
The direct product of finitely many algebraic groups over $\Gamma$ is again an algebraic group. Thus, given $a_{1},\ldots,a_{l}\in\Gamma\setminus\{1\}$ and $n_{1},\dots,n_{l},r\in{\mathbb{N}}$, the product $\Gamma^{r}\times Z(a_{1})^{n_{1}}\times\dots\times Z(a_{l})^{n_{l}}$ is also an algebraic group.
Example 1.4.
If $a$ is conjugate to $a^{\prime}=bab^{-1}$ then $V=Z(a)$ is isomorphic to $V^{\prime}=Z(a^{\prime})$ as an algebraic group, the isomorphism $v\mapsto bvb^{-1}$ and its inverse being algebraic maps. For this reason, if $\Gamma$ is a torsion-free hyperbolic group, any group as in Example 1.3 is isomorphic, as an algebraic group, to a group of the form $\Gamma^{r}\times\langle c_{1}\rangle^{n_{1}}\times\dots\times\langle c_{l}\rangle^{n_{l}}$ where $\langle c_{1}\rangle,\dots,\langle c_{l}\rangle$ are non-conjugate maximal cyclic groups.
Example 1.5.
[…]
$\begin{bmatrix}u_{1}\\ u_{2}\end{bmatrix}\circledast\begin{bmatrix}v_{1}\\ v_{2}\end{bmatrix}=\begin{bmatrix}u_{1}^{2}u_{2}^{-1}v_{1}^{2}v_{2}^{-1}u_{2}u_{1}^{-1}v_{2}v_{1}^{-1}\\ u_{2}u_{1}^{-1}v_{2}v_{1}^{-1}u_{1}^{2}u_{2}^{-1}v_{1}^{2}v_{2}^{-1}u_{2}u_{1}^{-1}v_{2}v_{1}^{-1}\end{bmatrix}$
but this is just an algebraic group isomorphic to $\Gamma^{2}$ with its standard multiplication law. Here we say that two algebraic groups $(V,\mu),(V^{\prime},\mu^{\prime})$ are isomorphic (as algebraic groups over $\Gamma$) if there is an isomorphism of algebraic varieties $F:V\rightarrow V^{\prime}$ which is a group isomorphism.
In this paper we work over a fixed torsion-free hyperbolic group $\Gamma$. A variety $V$ is called irreducible if it is not the union of two proper subvarieties. For an algebraic group, it is irreducible if and only if it is connected for the Zariski topology (see Section 4.3). We thus say that an algebraic group $V$ is connected if the underlying variety is connected, or equivalently irreducible. Focusing on connected algebraic groups is not a big restriction since any algebraic group has a finite index subgroup which is connected (this is a consequence of equational Noetherianity due to Sela [Sel09], see Section 4.3). All the examples of algebraic groups given above are connected.
Our main theorem gives a classification of all connected algebraic groups over $\Gamma$, and says that all of them come from the examples above.
Theorem 1.6.
Let $\Gamma$ be a torsion-free hyperbolic group. Let $V$ be a connected algebraic group over $\Gamma$.
Then there exist non-conjugate maximal cyclic subgroups $\langle c_{1}\rangle,\ldots,\langle c_{l}\rangle$ of $\Gamma$, and integers $r\in{\mathbb{N}}$ and $n_{1},\ldots,n_{l}\in{\mathbb{N}}\setminus\{0\}$ such that $V$ is isomorphic, as an algebraic group, to
$\Gamma^{r}\times\langle c_{1}\rangle^{n_{1}}\times\ldots\times\langle c_{l}\rangle^{n_{l}}$
where the multiplication law is the standard coordinate-wise product.
Moreover, this gives a complete classification of connected algebraic groups up to isomorphism: the conjugacy classes of the groups $\langle c_{1}\rangle,\dots,\langle c_{l}\rangle$ and the exponents $n_{i}$ and $r$ are uniquely determined (up to permutation) by the isomorphism class of $(V,\mu)$ as an algebraic group.
Remark 1.7.
As mentioned above, if $V$ is an algebraic group which is not assumed to be connected, then the connected component of $V$ containing the neutral element is a finite index subgroup which is itself a connected algebraic group. However, with the definition given above, this remark is not enough to deduce a description of all algebraic groups from our result on connected ones. Indeed, it is not clear even which finite groups can be realised as algebraic groups: every finite set is an algebraic set but its is not clear whether a given group law can be expressed as a word map.
With this in mind, it would be more natural to enlarge the category of algebraic groups by allowing the group law to be given by word maps only locally. For connected algebraic groups, this definition is equivalent to the previous one. However, it has the advantage of being stable under taking supergroups of finite index: with the new definition, if $G$ has a finite index subgroup which is isomorphic to a connected algebraic group, then $G$ itself is isomorphic to an algebraic group.
Other notions and questions
One can consider more general notions. A group $(V,\mu)$ is definable over a structure $\Gamma$ (such as a field, or a group) if $V$ is a subset of $\Gamma^{d}$ definable by a first order formula, and $\mu$ (or more precisely the graph of $\mu$) is also definable in this way. Even more generally, a group $(V,\mu)$ is interpretable over $\Gamma$ if $V=\tilde{V}/\sim$ where $\tilde{V}$ is a definable subset and $\sim$ is a definable equivalence relation, and similarly for the graph of $\mu$.
Over an algebraically closed field, the Weil-Hrushovski theorem states that any interpretable group is an algebraic group (see [Poi01] Theorem 4.13, with (3) from §4.1), where the notion of algebraic group here should not be restricted to affine algebraic groups but should also include abelian varieties for example.
[…]
Returning to the case where $\Gamma$ is a torsion-free hyperbolic group (and even a free group), we ask:
Problem 1.8.
Let $\Gamma$ be a torsion-free hyperbolic group. Classify the definable and interpretable groups over $\Gamma$.
We note that Byron and Sklinos proved, relying on Sela’s work on the first order theory of free groups, that no infinite field is interpretable over the free group [DBS19, Skl24, Sel06, Sel24].
Ingredients of the proof.
Our proof of Theorem 1.6 uses basic tools from algebraic geometry over groups [BMR99]. The central notion is that of the group of functions $L_{V}$ over a variety $V$ (also called coordinate group in the literature). For instance, in the case where $\Gamma$ is a free group, coordinate groups $L_{V}$ are precisely finitely generated residually free groups. When $V$ is irreducible (see below), $L_{V}$ is actually a limit group, a class of well behaved groups which has been widely studied [Sel01, KM98, Sel09, GW18]. The group $L_{V}$ contains a preferred copy of $\Gamma$ corresponding to the constant functions (except for the degenerate case $V=\emptyset$). Given a point $v$ in $V$, the evaluation homomorphism $L_{V}\rightarrow\Gamma$ is defined by evaluating a function $f\in L_{V}$ at the point $v$. This gives a bijective correspondence which identifies $V$ with the set $\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ of all homomorphisms $L_{V}\to\Gamma$ that restrict to the identity on the preferred copy of $\Gamma$ in $L_{V}$.
Of fundamental importance will be the functoriality of the correspondence between $V$ and $L_{V}$: a morphism of algebraic varieties $\alpha:V\to W$ corresponds to a group morphism $\alpha^{*}:L_{W}\to L_{V}$, and conversely. In particular, the group of automorphisms of $V$ (as an algebraic variety) is isomorphic to ${\mathrm{Aut}}_{\Gamma}(L_{V})$, the group of automorphisms of $L_{V}$ that are the identity on $\Gamma$. For example, if $V$ is homogeneous under the action of its automorphism group, then ${\mathrm{Aut}}_{\Gamma}(L_{V})$ acts transitively on $\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$. Using tools such as Grushko decompositions and JSJ decompositions, which help describe automorphisms of $L_{V}$, we get strong restrictions on $L_{V}$ and thus on $V$.
However this is not enough to complete the description of algebraic groups, since there are varieties $V$ that are homogeneous but do not correspond to algebraic groups over $\Gamma$. In Section 5.2 for instance, we describe a variety $V$ over $\Gamma$ a free group, on which there is a free transitive action of the mapping class group of a surface, but it follows from our classification that $V$ does not admit a structure of algebraic group.
If $V$ is an algebraic group, its action on itself yields a representation of $V$ into the group ${\mathrm{Aut}}_{\Gamma}(L_{V})$. The fact that this representation comes from an algebraic action of an algebraic group gives strong restrictions on the image of $V$ in ${\mathrm{Aut}}_{\Gamma}(L_{V})$. For example, in algebraic geometry over a field $k$, if an algebraic group acts algebraically on the affine space $k^{n}$, then any orbit for its action on the ring of functions $k[X_{1},\dots X_{n}]$ has bounded degree. The analogous statement here says that the action of $V$ on the group of functions $L_{V}$ has bounded stretch: there exists a constant $C$ such that all elements of $V$ act on $L_{V}$ with Lipschitz constant at most $C$ (see Lemma 7.5) where the length of elements of $L_{V}$ is measured in a word metric where all elements of $\Gamma\setminus\{1\}$ have length 1.
Because of the lack of local finiteness of the metric involved, this does not imply finiteness of the image of $V$. For instance, if $L_{V}=\Gamma*\langle X_{1},\dots,X_{n}\rangle$, and given two tuples $a=(a_{1},\dots,a_{n}),b=(b_{1},\dots,b_{n})\in\Gamma$, the automorphism $\tau_{a,b}:X_{i}\mapsto a_{i}X_{i}b_{i}$ is an automorphism whose powers have bounded Lipschitz constants, and which usually has infinite order. Still, this gives strong restrictions on the representation of $V$, and for instance, staying in the case where $L_{V}=\Gamma*\langle X_{1},\dots,X_{n}\rangle$, it prevents its image from containing complicated automorphisms such as fully irreducible automorphisms (in the sense of free products relative to $\Gamma$). Applying a result by Handel and Mosher [HM20] and generalized by Horbez and the first author [GH22], one deduces that there is a triangular structure invariant under the image of $V$ in ${\mathrm{Aut}}_{\Gamma}(L_{V})$ (the image of $V$ looks like a unipotent group).
The endgame is then algebraic: using the fact that the action of $V$ on $V$ is triangular, one can then express the associativity of the group law to deduce that our algebraic group is isomorphic to a standard one.
Organization of the paper
The paper is structured as follows: in Section 2, we survey tools of algebraic geometry over groups and the function group $L_{V}$ associated to a variety $V$.
In Section 3, we recall some results about $\Gamma$-limit groups, in particular we define the JSJ decomposition for such groups and give some of their properties.
In Section 4, we consider algebraic groups over $\Gamma$. We prove a useful formula expressing the group law under the identification of $V$ with ${\mathrm{Hom}}_{\Gamma}(L_{V})$, and express homogeneity of the underlying variety under this same identification.
We use this in Section 5 to deduce a number of constraints that the JSJ decomposition of the function group $L_{V}$ of a homogeneous irreducible variety must satisfy.
In Section 6, we construct a canonical algebraic subgroup $(W,\mu|_{W\times W})$ of $(V,\mu)$ associated to the abelian factors in $V$. We deduce that the JSJ decomposition of $L_{W}$ and of $L_{V}$ have no rigid vertices other than the one containing the constants group $\Gamma$.
In Section 7, we introduce the bounded stretch argument and use it to show that the JSJ decomposition of $L_{V}$ cannot contain any surface type vertices either.
This enables us in Section 8 to give a complete description of the various possibilities for the variety $V$, via a description of the JSJ decomposition of $L_{V}$.
Finally, in Section 9, we use the bounded stretch argument to find a triangular structure of the image of $V$ and $V^{2}$ in ${\mathrm{Aut}}_{\Gamma}(L_{V})$, and use it to find a change of coordinates under which the group law becomes the standard coordinate-wise product, thus completing the proof of Theorem 1.6.
Acknowledgement
The first author was partially supported by European Research Council (ERC GOAT 101053021), by the french ANR grant ANR-22-CE40-0004 GOFR, and benefited from the support of the French government ”Investissements d’Avenir” program integrated to France 2030 (ANR-11-LABX-0020-01). The second author was supported by the ISRAEL SCIENCE FOUNDATION (grant No. 2176/20) and CNRS.
2 Algebraic geometric tools over groups
In all this paper, we fix $\Gamma$ a non elementary torsion-free hyperbolic group.
2.1 $\Gamma$-groups.
Following [BMR99], a $\Gamma$-group $G$ is a group $G$ together with a specified monomorphism $\iota:\Gamma\rightarrow G$. A $\Gamma$-homomorphism $\varphi:G\rightarrow G^{\prime}$ is a homomorphism that commutes with the specified embeddings. We will usually keep $\iota$ implicit, and identify $\Gamma$ with its image; thus, a $\Gamma$-homomorphism is a homomorphism that restricts to the identity on $\Gamma$. We denote by ${\mathrm{Hom}}_{\Gamma}(G,G^{\prime})$ the set of $\Gamma$-homomorphisms.
For instance $\Gamma*\langle x_{1},\dots,x_{d}\rangle$ (where $\langle x_{1},\dots,x_{d}\rangle$ is the free group on the $x_{i}$’s) is a $\Gamma$-group whose elements are called $\Gamma$-words: they can be viewed as words in the variables $x_{1}^{\pm},\dots,x_{d}^{\pm}$ with coefficients in $\Gamma$.
A subset $S$ of a $\Gamma$-group $G$ is a $\Gamma$-generating set if $\Gamma\cup S$ generates $G$ as a group.
Definition 2.1.
A $\Gamma$-group $G$ is residually $\Gamma$ if for any $g\in G\setminus\{1\}$ there exists a $\Gamma$-homomorphism $h:G\to\Gamma$ such that $h(g)\neq 1$.
See Definition 2.11 for the stronger notion of being fully residually $\Gamma$.
Definition 2.2.
Given a $\Gamma$-group $G$, its largest residually $\Gamma$ quotient is the group
$G/K\quad\text{where }K=\bigcap_{h\in\mathrm{Hom}_{\Gamma}(G,\Gamma)}\ker h$
Clearly, $G/K$ is residually $\Gamma$ and if $G\to G^{\prime}$ is a $\Gamma$-homomorphism to a $\Gamma$-group $G^{\prime}$ which is residually $\Gamma$, it factors through $G/K$. If $G=\langle\Gamma,x_{1},\dots,x_{d}|\Sigma\rangle$, then $K$ is called the radical of $\Sigma$ in [BMR99].
2.2 Variety over $\Gamma$.
A system of equations over $\Gamma$ in the variables $x_{1},\ldots,x_{d}$ is a set of words
$\Sigma=\{w_{i}(x_{1},\ldots,x_{d})\mid i\in I\}\subset\Gamma*\langle x_{1},\ldots,x_{d}\rangle.$
The variety $V_{\Sigma}\subseteq\Gamma^{d}$ it defines is
$V_{\Sigma}=\{(v_{1},\ldots,v_{d})\in\Gamma^{d}\mid w_{i}(v_{1},\ldots,v_{d})=1\textrm{ for each }i\in I\}.$
Following Lemma 6(6) p. 49 and Proposition 1 p. 31 in [BMR99], since $\Gamma$ is CSA and non abelian, the collection of varieties is stable under finite union and is the collection of closed sets of a topology on $\Gamma^{d}$, called the Zariski topology (see Theorem 3 p. 50 in [BMR99]).
2.3 Group of functions.
Let $V$ be a subset of $\Gamma^{d}$. A word map on $V$ is a map $f:V\to\Gamma$ for which there exists a word $w(x_{1},\ldots,x_{d})\in\Gamma*\langle x_{1},\ldots,x_{d}\rangle$ such that for all $(v_{1},\dots,v_{d})\in V$, $f(v_{1},\ldots,v_{d})=w(v_{1},\ldots,v_{d})$.
For example, the projection on the $k$-th coordinate $V\to\Gamma,(v_{1},\ldots,v_{d})\mapsto v_{k}$ is a word map. We denote it by $X_{k}$.
Definition 2.3 (Group of functions).
The group of functions $L_{V}$ on $V$ is the set of word maps $f:V\to\Gamma$, endowed with multiplication in the range: given $f,f^{\prime}\in L_{V}$ we define $ff^{\prime}:V\to\Gamma$ by setting $ff^{\prime}(v)=f(v)f^{\prime}(v)$.
The group $L_{V}$ comes with the $\Gamma$-generating family $(X_{1},\dots,X_{d})$ formed by the coordinate functions.
By definition of a word map, there is a surjective map $\Gamma*\langle x_{1},\dots,x_{d}\rangle\rightarrow L_{V}$ sending a word $w$ to the corresponding word map, and sending the generator $x_{k}$ to the coordinate map $X_{k}$. Thus, $(X_{1},\dots,X_{d})$ is indeed a $\Gamma$-generating family of $L_{V}$.
The group of functions on $V$ is also known as its coordinate group [BMR99].
All the varieties we consider will be non empty. In this case, $L_{V}$ is a $\Gamma$-group where $\Gamma$ embeds in $L_{V}$ as the set of constant maps.
2.4 Equations of the variety and functions on $V$.
If $V$ is a variety defined by $\Sigma=\{w_{i}(x_{1},\dots,x_{d})\mid i\in I\}$, then the map $\Gamma*\langle x_{1},\dots,x_{d}\rangle\rightarrow L_{V}$ vanishes on each $w_{i}$, hence factors through the group $L_{\Sigma}=\langle\Gamma,x_{1},\dots,x_{d}\mid w_{i},i\in I\rangle$.
Given a point $v=(v_{1},\dots,v_{d})\in V$, we denote the evaluation homomorphism by $h_{v}:\Gamma*\langle x_{1},\dots,x_{d}\rangle\rightarrow\Gamma$, it is defined by $h_{v}(w)=w(v_{1},\dots,v_{d})$ for any word $w\in\Gamma*\langle x_{1},\dots,x_{d}\rangle$. This is a $\Gamma$-homomorphism so $h_{v}\in{\mathrm{Hom}}_{\Gamma}(\Gamma*\langle x_{1},\dots,x_{d}\rangle,\Gamma)$.
Every evaluation homomorphism $h_{v}$ factors through $L_{\Sigma}$ and through $L_{V}$. We still call the factored map $L_{\Sigma}\rightarrow\Gamma$ or $L_{V}\rightarrow\Gamma$ an evaluation homomorphism and denote it by $h_{v}$.
Lemma 2.4.
Every $h\in{\mathrm{Hom}}_{\Gamma}(L_{\Sigma},\Gamma)$ agrees with some evaluation homomorphism $h_{v}:L_{\Sigma}\to\Gamma$.
Proof.
Given $h\in{\mathrm{Hom}}_{\Gamma}(L_{\Sigma},\Gamma)$, consider the point $v_{h}=(h(\bar{x}_{1}),\dots,h(\bar{x}_{d}))\in\Gamma^{d}$ where we denote by $\bar{x}_{i}\in L_{\Sigma}$ the image of $x_{i}$ under the morphism $\Gamma*\langle x_{1},\dots,x_{d}\rangle\rightarrow L_{\Sigma}$. The point $v_{h}$ lies in the variety $V$. Indeed, for each $w_{i}\in\Sigma$, we have $w_{i}(h(\bar{x}_{1}),\dots,h(\bar{x}_{d}))=h(w_{i}(\bar{x}_{1},\dots,\bar{x}_{d}))=1$ because $h$ is a $\Gamma$-morphism. We check that $h=h_{v_{h}}$: given any word $w\in\Gamma*\langle x_{1},\dots,x_{d}\rangle$ with image $\bar{w}\in L_{\Sigma}$, we have
$h_{v_{h}}(\bar{w})=w(h(\bar{x}_{1}),\dots,h(\bar{x}_{d}))=h(w(\bar{x}_{1},\ldots,\bar{x}_{d}))=h(\bar{w}).$
This concludes the proof. ∎
Two words in $\Gamma*\langle x_{1},\dots,x_{d}\rangle$ represent the same function on $V$ if and only if they have the same image under all evaluations maps $h_{v}$ for $v\in V$. Thus one can alternatively describe $L_{V}$ as the quotient
$L_{V}=L_{\Sigma}/\bigcap_{v\in V}\ker h_{v}.$
By construction, $L_{V}$ is residually $\Gamma$: for every function $f\in L_{V}\setminus\{1\}$ there is a $\Gamma$-morphism $L_{V}\rightarrow\Gamma$ which does not kill $f$, namely the evaluation map $h_{v}$ at a point $v\in V$ where $f$ does not vanish.
As an immediate consequence of Lemma 2.4 we get
Corollary 2.5.
If $V$ the variety defined by the set of equations $\Sigma$, then $L_{V}$ is the largest quotient of $L_{\Sigma}$ that is residually $\Gamma$.
In particular, the canonical map from ${\mathrm{Hom}}_{\Gamma}(L_{V},\Gamma)$ to ${\mathrm{Hom}}_{\Gamma}(L_{\Sigma},\Gamma)$ is a bijection.
Example 2.6.
Fix an element $c\in\Gamma\setminus\{1\}$ which is not a proper power so that $Z(c)=\langle c\rangle$. Let $V=Z(c)^{d}\subset\Gamma^{d}$ be the variety defined by the system of equations $\Sigma=\{[x_{1},c],\dots,[x_{d},c]\}$. Here $L_{\Sigma}=\langle\Gamma,x_{1},\dots,x_{d}\mid[x_{1},c],\ldots,[x_{d},c]\rangle$ and
$L_{V}\simeq\Gamma*_{\langle c\rangle}(\langle c\rangle\oplus{\mathbb{Z}}^{d})$
is its largest residually $\Gamma$ quotient (see for instance [BMR02, Theorem 1] to see it is indeed residually $\Gamma$).
2.5 Correspondence $V\leftrightarrow{\mathrm{Hom}}_{\Gamma}(L_{V},\Gamma)$.
Consider the maps
$\left\{\begin{matrix}V&\rightarrow&{\mathrm{Hom}}_{\Gamma}(L_{V},\Gamma)\\ v&\mapsto&h_{v}\end{matrix}\right.\text{\quad and \quad}\left\{\begin{matrix}{\mathrm{Hom}}_{\Gamma}(L_{V},\Gamma)&\rightarrow&V\\ h&\mapsto&v_{h}=(h(X_{1}),\dots,h(X_{d}))\end{matrix}\right.$
The first associates to a point of $V$ its evaluation map, the second is the map $h\mapsto v_{h}$ that appears in the proof of the previous lemma, up to the bijection ${\mathrm{Hom}}_{\Gamma}(L_{V},\Gamma)\rightarrow{\mathrm{Hom}}_{\Gamma}(L_{\Sigma},\Gamma)$ mentioned in Corollary 2.5.
We claim that these maps are inverses of each other. Indeed, we already proved in Lemma 2.4 that $h_{v_{h}}=h$ for all $h\in{\mathrm{Hom}}_{\Gamma}(L_{\Sigma},\Gamma)$. This is true on ${\mathrm{Hom}}_{\Gamma}(L_{V},\Gamma)$ via the canonical bijection between ${\mathrm{Hom}}_{\Gamma}(L_{\Sigma},\Gamma)$ and ${\mathrm{Hom}}_{\Gamma}(L_{V},\Gamma)$. For the other direction, we consider $v\in V$ and $h_{v}:L_{V}\rightarrow\Gamma$ the corresponding evaluation morphism. We have
$v_{h_{v}}=(h_{v}(X_{1}),\dots,h_{v}(X_{d}))=(X_{1}(v),\dots,X_{d}(v))=v.$
2.6 Algebraic maps.
Let $V\subseteq\Gamma^{d},V^{\prime}\subseteq\Gamma^{d^{\prime}}$ be algebraic varieties over $\Gamma$. An algebraic map from $V$ to $V^{\prime}$ is a map $F:V\to V^{\prime}$ such that there exist word maps $f_{1},\dots,f_{d^{\prime}}:V\rightarrow\Gamma$ such that $F(v)=(f_{1}(v),\ldots,f_{d^{\prime}}(v))$ for all $v\in V$. An algebraic automorphism $F:V\rightarrow V$ is a bijective algebraic map whose inverse is an algebraic map.
Recalling that elements of $L_{V}$ (respectively $L_{V^{\prime}}$) are word maps $V\to\Gamma$ (respectively $V^{\prime}\to\Gamma$) any algebraic map $F:V\rightarrow V^{\prime}$ induces a $\Gamma$-homomorphism
$F^{*}:\left\{\begin{matrix}L_{V^{\prime}}&\to&L_{V}\\ f&\mapsto&f\circ F.\end{matrix}\right.$
Note that given a variety $V^{\prime\prime}$ and an algebraic map $G:V^{\prime}\to V^{\prime\prime}$, we have $(G\circ F)^{*}=F^{*}\circ G^{*}$. In other words, there is a contravariant functor from the category of algebraic varieties over $\Gamma$ to the category of $\Gamma$-groups that are residually $\Gamma$.
Conversely, given a $\Gamma$-homomorphism $\theta:L_{V^{\prime}}\to L_{V}$ (with $\Gamma$-generating families $(X_{1},\dots,X_{d})$ and $(X^{\prime}_{1},\dots,X^{\prime}_{d^{\prime}})$ of $L_{V}$ and $L_{V^{\prime}}$ respectively) there is an induced map $\theta_{*}:{\mathrm{Hom}}_{\Gamma}(L_{V},\Gamma)\to{\mathrm{Hom}}_{\Gamma}(L_{V^{\prime}},\Gamma)$ given by $h\mapsto h\circ\theta$. Via the identification between $V$ and ${\mathrm{Hom}}_{\Gamma}(L_{V},\Gamma)$ (respectively between $V^{\prime}$ and ${\mathrm{Hom}}_{\Gamma}(L_{V^{\prime}},\Gamma)$) induced by the $\Gamma$-generating families, this can be seen as an algebraic map $V\to V^{\prime}$: the image of $X^{\prime}_{1},\dots,X^{\prime}_{d^{\prime}}\in L_{V^{\prime}}$ under $\theta$ can be written using the $\Gamma$-generating set $\{X_{1},\dots,X_{d}\}$ of $L_{V}$, say $\theta(X^{\prime}_{i})=u_{i}(X_{1},\dots,X_{d})$ for some $\Gamma$-words $u_{i}\in\Gamma*\langle x_{1},\dots,x_{d}\rangle$, and the algebraic map $V\rightarrow V^{\prime}$ is defined by
$(v_{1},\dots,v_{d})\in V\mapsto(u_{1}(v_{1},\dots,v_{d}),\dots,u_{d^{\prime}}(v_{1},\dots,v_{d})).$
In particular, there is an isomorphism between the group ${\mathrm{Aut}}(V)$ of algebraic automorphisms of the variety $V$, and the group ${\mathrm{Aut}}_{\Gamma}(L_{V})$ of $\Gamma$-automorphisms of the $\Gamma$-group $L_{V}$.
2.7 Noetherianity and irreducible varieties.
Sela proved the following
Theorem 2.7 ([Sel09]).
Torsion-free hyperbolic groups are equationally Noetherian: for any system of equations $\Sigma\subset\Gamma*\langle x_{1},\dots,x_{n}\rangle$, there exists a finite subsystem $\Sigma^{\prime}\subset\Sigma$ such that $V_{\Sigma}=V_{\Sigma^{\prime}}$.
Equivalently, any descending chain of varieties $V_{1}\supset V_{2}\supset V_{3}\ldots$ is stationary.
Definition 2.8.
A non-empty variety is irreducible if it cannot be written as a finite union of proper subvarieties.
As a straightforward corollary of Theorem 2.7 we get that any variety $V$ can be written as a finite union $V_{1}\cup\ldots\cup V_{s}$ of irreducible varieties. Additionally, there is a unique such decomposition for which $s$ is minimal (equivalently, no inclusion $V_{i}\subset V_{j}$ holds for $i\neq j$): in this case, $V_{1},\dots,V_{s}$ are called the irreducible components of $V$.
Definition 2.9.
Let $G,G^{\prime}$ be two $\Gamma$-groups. A collection ${\mathcal{H}}$ of $\Gamma$-homomorphisms $G\rightarrow G^{\prime}$ is discriminating if for every finite subset $F\subset G$ there exists $h\in{\mathcal{H}}$ such that $h$ is injective in restriction to $F$.
Remark 2.10.
For a sequence of $\Gamma$-homomorphisms, we will use the following variant of this definition: a sequence of $\Gamma$-homomorphisms $h_{i}:G\rightarrow G^{\prime}$ ($i\in{\mathbb{N}}$) is discriminating if for every finite subset $F\subset G$, for all $i$ large enough, $h_{i}$ is injective in restriction to $F$. Note that for $G$ countable, a collection ${\mathcal{H}}$ of $\Gamma$-homomorphisms is discriminating if and only if it contains a discriminating sequence.
Equivalently, ${\mathcal{H}}$ is discriminating if for any finite collection of non-trivial elements $g_{1},\dots,g_{n}\in G\setminus\{1\}$, there exists $h\in{\mathcal{H}}$ such that $h(g_{1}),\dots,h(g_{n})$ are all non-trivial.
Definition 2.11.
A $\Gamma$-group $G$ is fully residually $\Gamma$ if it admits a discriminating family of $\Gamma$-homomorphisms to $\Gamma$ (equivalently ${\mathrm{Hom}}_{\Gamma}(G,\Gamma)$ is discriminating).
Sela proved that the class of finitely generated $\Gamma$-groups which are fully residually $\Gamma$ coincide with that of $\Gamma$-limit groups [Sel09] (actually, in [Sel09] §1, Sela works with equations which do not involve constants in $\Gamma$, so the $\Gamma$-limit groups there don’t come with a preferred copy $\Gamma$).
The following lemma is a slight variation on [BMR99, Theorem D2].
Lemma 2.12.
The variety $V$ is irreducible if and only if $L_{V}$ is fully residually $\Gamma$.
In this case, if $V_{0}\subset V$ is a dense subset (for the Zariski topology), then the family of evaluation homomorphisms $\{h_{v}|v\in V_{0}\}\subset\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ is discriminating on $L_{V}$.
Proof.
We first prove the second assertion by contradiction: we assume that $V$ is irreducible, but that $\{h_{v}|v\in V_{0}\}$ is not discriminating. Then there exists $f_{1},\dots,f_{n}\in L_{V}\setminus\{1\}$ such that for each $v\in V_{0}$, there exists $i\leq n$ such that $h_{v}(f_{i})=f_{i}(v)=1$. Denote by $V_{i}=f_{i}^{-1}(\{1\})$ the subvariety of $V$ defined by $f_{i}$. In other words, $V_{0}\subset\bigcup_{i}V_{i}$. Since $V_{0}$ is dense and the $V_{i}$ are closed, we get $V=\bigcup_{i}V_{i}$ so by irreducibility, $V=V_{i}$ for some $i\leq n$. This means that $f_{i}$ vanishes on $V$, i. e. $f_{i}=1$ a contradiction.
This proves the second assertion. The direct implication of the first assertion follows.
Conversely, assume that $V=V_{1}\cup\dots\cup V_{n}$ is a union of proper subvarieties. For each $i\leq n$, consider a word map $f_{i}\in L_{V}\setminus\{1\}$ that vanishes on $V_{i}$. Then for any point $v\in V$, $v$ lies in $V_{i}$ for some $i\leq n$ and the evaluation homomorphism $h_{v}$ vanishes on $f_{i}$. This shows that the collection of evaluation homomorphisms is not discriminating. Since by Lemma 2.4 every $\Gamma$-morphism to $\Gamma$ is an evaluation homomorphism, this proves the lemma. ∎
2.8 Direct products
Let $V_{1},V_{2}$ be two non-empty algebraic varieties, and let $V=V_{1}\times V_{2}$ be the product variety. We now describe the group of functions $L_{V}$ in terms of $L_{V_{1}}$ and $L_{V_{2}}$.
For $i\leq 2$, denote by $p_{i}:V\rightarrow V_{i}$ the $i$-th projection. Recall that the map $p_{i}^{*}:L_{V_{i}}\rightarrow L_{V}$ sends a word map $f:V_{i}\rightarrow\Gamma$ to $f\circ p_{i}$. The two maps $p_{1}^{*},p_{2}^{*}$ are injective, they agree on $\Gamma$, and thus yield a $\Gamma$-homomorphism $\varphi:L_{V_{1}}*_{\Gamma}L_{V_{2}}\rightarrow L_{V}$.
Lemma 2.13.
$L_{V}$ is the maximal residually $\Gamma$ quotient of $L_{V_{1}}*_{\Gamma}L_{V_{2}}$.
More precisely, the map $L_{V_{1}}*_{\Gamma}L_{V_{2}}\rightarrow L_{V}$ is surjective and its kernel is exactly the intersection of all kernels of $\Gamma$-homomorphism $L_{V_{1}}*_{\Gamma}L_{V_{2}}\rightarrow\Gamma$.
Proof.
Let $\Sigma_{1}\subset\Gamma*\langle x_{1},\dots,x_{d}\rangle$ (resp. $\Sigma_{2}\subset\Gamma*\langle y_{1},\dots,y_{d^{\prime}}\rangle$) be a system of equations defining $V_{1}$ (resp. $V_{2}$). By Lemma 2.4, $L_{V_{1}}$ is the maximal residually-$\Gamma$ quotient of the group
$L_{\Sigma_{1}}=\langle\Gamma,x_{1},\dots,x_{d}|\Sigma_{1}\rangle$
(i. e. its quotient by the intersection $K_{1}$ of all $\Gamma$-homomorphisms $L_{\Sigma_{1}}\to\Gamma$), and similarly for $L_{V_{2}}$. Since $\Sigma_{1}\cup\Sigma_{2}\subset\Gamma*\langle x_{1},\dots,x_{d},y_{1},\dots,y_{d^{\prime}}\rangle$ is a system of equations defining $V$, $L_{V}$ is the maximal residually $\Gamma$ quotient of
$L_{\Sigma_{1}}*_{\Gamma}L_{\Sigma_{2}}=\langle\Gamma,x_{1},\dots,x_{d},y_{1},\dots,y_{d^{\prime}}|\Sigma_{1}\cup\Sigma_{2}\rangle.$
Clearly $K_{1}$ and $K_{2}$ die in $L_{V}$ so $L_{V}$ is also the maximal residually $\Gamma$ quotient of $L_{V_{1}}*_{\Gamma}L_{V_{2}}$.
Remark 2.14.
One can prove that if $V_{1}$ and $V_{2}$ are irreducible, then so is $V_{1}\times V_{2}$. Indeed, the proof over fields given for instance in [Har83, Exercise I.3.15] adapts directly.
Example 2.15.
The description given by the lemma is not very explicit. We will need a more precise description only in simple situations such as the following. Start with $V_{1}=V_{2}=Z(c)$ where $c$ is not a proper power. Then for $i\in\{1,2\}$,
$L_{V_{i}}=\Gamma*_{\langle c\rangle}(\langle c\rangle\oplus\langle t_{i}\rangle)\simeq\Gamma*_{\langle c\rangle}{\mathbb{Z}}^{2}$
and
$L_{V_{1}\times V_{2}}=\Gamma*_{\langle c\rangle}(\langle c\rangle\oplus\langle t_{1}\rangle\oplus\langle t_{2}\rangle).$
The map $L_{V_{1}}*_{\Gamma}L_{V_{2}}\rightarrow L_{V_{1}\times V_{2}}$ is not injective: its kernel contains for example the commutator $[t_{1},t_{2}]$.
3 $\Gamma$-limit groups
3.1 Description of the JSJ decomposition of $\Gamma$-limit groups
In this section we follow the terminology of [GL17].
Let $L$ be a $\Gamma$-limit group. Given two collections ${\mathcal{A}},{\mathcal{H}}$ of subgroups of $L$, an $({\mathcal{A}},{\mathcal{H}})$-tree is a tree with a minimal action of $L$ such that each edge stabilizer lies in ${\mathcal{A}}$, and each group in ${\mathcal{H}}$ is elliptic. A subgroup $H\subset L$ is $({\mathcal{A}},{\mathcal{H}})$-universally elliptic if it is elliptic in every $({\mathcal{A}},{\mathcal{H}})$-tree. An $({\mathcal{A}},{\mathcal{H}})$-JSJ decomposition is an $({\mathcal{A}},{\mathcal{H}})$-tree whose edge groups are $({\mathcal{A}},{\mathcal{H}})$-universally ellipic, and maximal for domination (given two $({\mathcal{A}},{\mathcal{H}})$-trees $T$, $T^{\prime}$, we say that $T$ dominates $T^{\prime}$ if vertex stabilizers of $T$ are elliptic in $T^{\prime}$).
In what follows, we will take for ${\mathcal{A}}$ the collection of abelian subgroups of $L$. Denoting by ${\mathcal{A}}_{nc}\subset{\mathcal{A}}$ the subcollection of all non-cyclic abelian groups, we will take =${\mathcal{H}}=\{\Gamma\}\cup{\mathcal{A}}_{nc}$. Thus $({\mathcal{A}},{\mathcal{H}})$-trees are abelian splittings of $L$ relative to $\Gamma$ and to non-cyclic abelian groups.
If $L$ is freely indecomposable relative to $\Gamma$, it has a canonical JSJ tree over abelian groups relative to $\Gamma$ and to non-cyclic abelian groups. Indeed $L$ is torsion-free and CSA (because $\Gamma$ is) thus Theorem 9.5 of [GL17] implies that the JSJ deformation space of $L$ relative to $\Gamma$ and to non-cyclic abelian subgroups exists. Its tree of cylinders is itself a JSJ tree, and is invariant by ${\mathrm{Aut}}_{\Gamma}(L)$. We denote by $\Lambda$ the corresponding graph of groups. We call the decomposition $L=\pi_{1}(\Lambda)$ the canonical JSJ decomposition of $L$. We will denote by $v_{\Gamma}$ the image in $\Lambda$ of the unique vertex $\tilde{v}_{\Gamma}$ of the JSJ tree fixed by $\Gamma$. We denote by $G_{v_{\Gamma}}=G_{\tilde{v}_{\Gamma}}$ the corresponding vertex group (which contains $\Gamma$).
If we drop the assumption of free indecomposability, we choose $L=L_{0}*\ldots*L_{r}*{\mathbb{F}}_{s}$ a Grushko decomposition relative to $\Gamma$ with $\Gamma\subset L_{0}$.
As mentioned above, we can consider the canonical JSJ-decomposition $L_{0}=\pi_{1}(\Lambda_{0})$ relative to $\Gamma$ and non-cyclic abelian groups. For each $i\geq 1$, $L_{i}$ has no preferred copy of $\Gamma$ (it contains no conjugate of $\Gamma$). There is an ${\mathrm{Aut}}(L_{i})$-invariant JSJ-decomposition $L_{i}=\pi_{1}(\Lambda_{i})$ relative to non-cyclic abelian groups.
The following definition will be convenient.
Definition 3.1 (Standard JSJ decomposition).
Gluing a wedge of $s$ edges with trivial edge group on $v_{\Gamma}$, and connecting $v_{\Gamma}$ to each graph of groups $\Lambda_{i}$ ($i>0$) using edges with trivial edge group (see Figure 1, one obtains a splitting of $L$. We call such a splitting a standard JSJ decomposition of $L$.
Figure 1: A typical standard JSJ decomposition. Blue edges have trivial edge group, black edges have non-trivial abelian edge group.
A standard JSJ decomposition of $L$ is indeed a JSJ decomposition of $L$ over abelian groups relative to $\Gamma$ and to non-cyclic abelian groups (but it is not invariant under ${\mathrm{Aut}}_{\Gamma}(L)$ in general).
Trees of cylinders can be characterized in the following way:
Lemma 3.2.
Let $G=\pi_{1}(\Lambda)$ be an abelian splitting of a CSA group $G$ without trivial edge groups.
The tree corresponding to $\Lambda$ is its own tree of cylinders if and only if it is bipartite between vertices with abelian and non-abelian stabilizer, and for every vertex $v$ with non-abelian stabilizer,
  • for every edge $e$ incident on $v$, the group $G_{e}$ is maximal abelian in $G_{v}$;
  • for every pair of edges $e\neq e^{\prime}$ incident on $v$, $G_{e}$ is not $G_{v}$-conjugate to $G_{e^{\prime}}$.
Proof.
In this context, cylinders have maximal abelian stabilizers, thus it is easy to see that a tree of cylinder will satisfy the conclusions of the lemma.
For the other direction, note that the given conditions imply that cylinders are exactly the set of edges adjacent to a common abelian vertex. Applying the tree of cylinders construction thus yields an isomorphic tree. ∎
We gather some properties of the standard JSJ decomposition which we will use.
A vertex $v$ in an abelian splitting $L=\pi_{1}(\Lambda)$ is rigid if $G_{v}$ is elliptic in every abelian splitting of $L$ relative to $\Gamma$ and to non-cyclic abelian groups. It is flexible otherwise.
The vertex $v$ is of abelian type if its vertex group $G_{v}$ is abelian. It is of surface type (see QH with trivial fiber in [Sel97, GL17]) if $G_{v}$ can be written as the fundamental group of a surface $\Sigma$ so that each incident edge group is conjugate to a subgroup of the fundamental group of a boundary component of $\Sigma$ (orbifolds are allowed in [GL17], but they don’t occur here because the groups we consider are torsion-free).
Lemma 3.3.
Let $L$ be a $\Gamma$-group which is also a $\Gamma$-limit group. A standard JSJ decomposition $\Lambda$ for $L$ satisfies the following
  1. 1.
    non-abelian flexible vertex groups are of surface type;
  2. 2.
    if $H\subset L$ is $({\mathcal{A}},{\mathcal{H}})$-universally elliptic and is contained in a surface type vertex group of $\Lambda$, it is contained in a boundary component;
  3. 3.
    the vertex group containing $\Gamma$ is rigid;
  4. 4.
    any edge $e$ with nontrivial $G_{e}$ joins a vertex with abelian group to a vertex with non abelian group;
  5. 5.
    for any non-abelian vertex $v$, and incident edge $e$, if $G_{e}$ is nontrivial it is maximal abelian (and thus malnormal) in $G_{v}$;
  6. 6.
    for any non-abelian vertex $v$, and every pair of edges $e\neq e^{\prime}$ with non-trivial stabilizer incident on $v$, $G_{e}$ is not $G_{v}$-conjugate to $G_{e^{\prime}}$.
Proof.
Item 1 is given by Theorem 9.5 of [GL17]. Item 2 appears as point (3) of Theorem 9.4 of [GL17], and implies item 3 by considering the case where $H=\Gamma$ (recall that we assume that $\Gamma$ is a non-elementary hyperbolic group). Items 4, 5 and 6 follow from Lemma 3.2. ∎
The following result, which follows from Proposition 4.15 of [GL17] characterizes (maybe not standard) JSJ trees among all $({\cal A},{\mathcal{H}})$-tree.
Lemma 3.4.
Let $\Delta$ be a splitting of $L$ corresponding to an $({\cal A},\Gamma)$-tree. Suppose the following hold:
  • edge groups of $\Delta$ are $({\mathcal{A}},{\mathcal{H}})$-universally elliptic;
  • a vertex group $G_{v}$ of $\Delta$ is either of surface type or rigid relative to the incident edge groups $G_{e}$ (that is, does not admit any non trivial $({\cal A},{\mathcal{H}}_{|G_{v}}\cup\{G_{e}\}_{o(e)=v})$-splitting, where ${\mathcal{H}}_{|G_{v}}$ is the set of conjugates of $H$ contained in $G_{v}$).
Then $\Delta$ is a JSJ decomposition for $L$.
4 Basic facts about algebraic groups
4.1 Algebraic groups over $\Gamma$.
Note that if $V\subset\Gamma^{d}$ is a variety, then $V\times V$ is a variety in $\Gamma^{2d}$.
Definition 4.1.
An algebraic group $(V,\mu)$ over $\Gamma$ is an algebraic variety $V$ together with a group law $\mu:V\times V\rightarrow V$ which is an algebraic map.
Note that we could have required in addition that the inverse map $V\rightarrow V$, defined by $g\mapsto g^{-1}$, be an algebraic map. In fact, when $\Gamma$ is torsion-free hyperbolic and $V$ is irreducible, our classification result shows that the inverse map is automatically algebraic.
[…]
For any $v,w\in V$ we have by associativity $\lambda_{v}\circ\lambda_{w}=\lambda_{v\circledast w}$ but $\rho_{v}\circ\rho_{w}=\rho_{w\circledast v}$. We also note that $\rho_{\check{v}}=\rho_{v}^{-1}$ and $\lambda_{\check{v}}=\lambda^{-1}_{v}$.
This induces automorphisms
$\left\{\begin{array}[]{rcl}\rho_{v}^{*}:L_{V}&\rightarrow&L_{V}\\ f&\mapsto&f\circ\rho_{v}\end{array}\right.\text{\quad and \quad}\left\{\begin{array}[]{rcl}\lambda_{v}^{*}:L_{V}&\rightarrow&L_{V}\\ f&\mapsto&f\circ\lambda_{v}\end{array}\right..$
and therefore gives two representations of the algebraic group $V$ into ${\mathrm{Aut}}_{\Gamma}(L_{V})$ by
$\left\{\begin{array}[]{rcl}V&\rightarrow&{\mathrm{Aut}}_{\Gamma}(L_{V})\\ v&\mapsto&\rho^{*}_{v}\end{array}\right.\text{\quad and \quad}\left\{\begin{array}[]{rcl}V&\rightarrow&{\mathrm{Aut}}_{\Gamma}(L_{V})\\ v&\mapsto&(\lambda^{*}_{v})^{-1}\end{array}\right.$
4.2 The group law formula
Let $V$ be an algebraic group. Any point $v\in V$ corresponds to a $\Gamma$-homomorphism $h_{v}\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$. We wish to understand the group law when points of $V$ are interpreted as elements of $\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$. We denote by $e$ the identity element of $V$ and by $h_{e}\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ the corresponding $\Gamma$-homomorphism.
Lemma 4.2.
For all $v,v^{\prime}\in V$, one has
$h_{v\circledast v^{\prime}}=h_{v}\circ\rho_{v^{\prime}}^{*}.$
In particular, for all $v\in V$, one has $h_{v}=h_{e}\circ\rho_{v}^{*}$.
Proof.
For any $f\in L_{V}$ (viewed as a function on $V$), $h_{v}\circ\rho_{v^{\prime}}^{*}(f)=h_{v}(f\circ\rho_{v^{\prime}})=f\circ\rho_{v^{\prime}}(v)=f(v\circledast v^{\prime})=h_{v\circledast v^{\prime}}(f)$. ∎
As an immediate consequence of the two equalities in Lemma 4.2 we get
Corollary 4.3.
For all $v,v^{\prime}\in V$ we have
$h_{v\circledast v^{\prime}}=h_{v}\circ\rho_{v^{\prime}}^{*}=h_{e}\circ\rho_{v}^{*}\circ\rho_{v^{\prime}}^{*}$
The following is a straightforward but useful consequence of the transitivity of the action of $G$ on itself by right multiplication.
Lemma 4.4.
If $V$ is an algebraic group, then for all $h,h^{\prime}\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$, there exists $v\in V$ such that $h^{\prime}=h\circ\rho_{v}^{*}$.
In particular, if $V$ is a connected algebraic group, then for any $h\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ the family of $\Gamma$-homomorphisms $\{h\circ\rho_{v}^{*}\mid v\in V\}$ is discriminating.
Proof.
Using the correspondence between $V$ and $\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$, the transitivity of the right action of $V$ on itself translates into the first assertion.
The second part follows from the fact that the family of all morphisms in $\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ is discriminating because $L_{V}$ is a $\Gamma$-limit group. ∎
4.3 Connected components of an algebraic group.
Let $(V,\mu)$ be an algebraic group over $\Gamma$. The following easy remarks are completely standard.
Let $V=V_{1}\cup\ldots\cup V_{s}$ be the unique (minimal) decomposition of $V$ as a finite union of irreducible varieties. Then the irreducible components are disjoint. Indeed, as a consequence of homogeneity all the points in $V$ are contained in the same number $k$ of irreducible components $V_{j}$. If $k>1$ we have $V_{1}=(V_{1}\cap V_{2})\cup\ldots\cup(V_{1}\cap V_{s})$, which contradicts the irreducibility of $V_{1}$.
It follows that the decomposition of $V$ into irreducible components is also its decomposition into connected components for the Zariski topology.
The connected component containing $e$ is a finite index algebraic subgroup of $V$.
Lemma 4.5.
Let $V$ be a connected algebraic group, and let $V_{0}\subset V$ be a finite index (not necessarily algebraic) subgroup of $V$. Then $V_{0}$ is dense in $V$.
Proof.
If $W$ is a subvariety of $V$ containing $V_{0}$, then finitely many translates of $W$ cover $V$. Since $V$ is irreducible, $W=V$. ∎
5 Homogeneous varieties
Algebraic groups are in particular homogeneous varieties. In this section, we collect a few consequences of homogeneity.
Definition 5.1.
A variety $V$ is homogeneous if the group of algebraic automorphisms $V\rightarrow V$ acts transitively on $V$.
Question 5.2.
What can we say about homogeneous varieties in general? Can we classify them ?
5.1 Homogeneity
The following is the reformulation of the definition of homogeneity using the correspondence $V\leftrightarrow{\mathrm{Hom}}_{\Gamma}(L_{V},\Gamma)$.
Lemma 5.3.
$V$ is homogeneous if and only if for all $h,h^{\prime}\in{\mathrm{Hom}}(L_{V},\Gamma)$, there exists $\alpha\in{\mathrm{Aut}}_{\Gamma}(L_{V})$ with $h\circ\alpha=h^{\prime}$.∎
We saw in Lemma 2.12 that $V$ is irreducible as an algebraic variety if and only if the family ${\mathrm{Hom}}_{\Gamma}(L_{V},\Gamma)$ is discriminating. Together with the previous lemma this gives
Corollary 5.4.
If $V$ is a homogeneous irreducible variety, for every $h\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$, the family $\{h\circ\alpha\mid\alpha\in{\mathrm{Aut}}_{\Gamma}(L_{V})\}$ is discriminating.∎
Note that Lemma 4.4 gives a more precise version of this result in the case where $V$ is an algebraic group.
The following lemma says in particular that Corollary 5.4 still holds if one replaces ${\mathrm{Aut}}_{\Gamma}(L_{V})$ by a finite index subgroup.
Lemma 5.5.
Let $h\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ and let $M$ be a subgroup of ${\mathrm{Aut}}_{\Gamma}(L_{V})$. Assume that $\{h\circ\alpha\mid\alpha\in M\}$ is a discriminating family.
Then for any finite index subgroup $M_{0}$ of $M$, the subfamily $\{h\circ\alpha\mid\alpha\in M_{0}\}$ is still discriminating.
Proof.
Write $M=M_{0}\varphi_{1}\sqcup\dots\sqcup M_{0}\varphi_{n}$ for some $\varphi_{1},\dots,\varphi_{n}\in M$. Fix a finite set $\{g_{1},\ldots,g_{r}\}\subseteq L_{V}$. Since $\{h\circ\alpha|\alpha\in M\}$ is discriminating, there exists $\alpha\in M$ such that $h\circ\alpha$ does not kill any $\varphi^{-1}_{i}(g_{j})$. Let $k$ be such that $\alpha=\alpha_{0}\varphi_{k}$ for some $\alpha_{0}\in M_{0}$. Then $h\circ\alpha_{0}$ does not kill any $g_{j}$. ∎
5.2 An example of homogeneous variety
Let $\Gamma$ be a torsion-free hyperbolic group and let $c\in\Gamma\setminus\{1\}$ be such that $\langle c\rangle$ is a maximal cyclic group of $\Gamma$ and such that any endomorphism of $\Gamma$ fixing $c$ is an automorphism (such a $c$ is known as a test element in $\Gamma$). Then the double $L_{V}=\Gamma*_{c}\bar{\Gamma}$ is a $\Gamma$-limit group, and it is the function group of a homogeneous variety: for any $h\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ we have that $h_{|\bar{\Gamma}}$ sends $\bar{c}$ to $c$, so it must be an isomorphism. Therefore $h=r\circ\sigma$ where $r$ is the canonical retraction and $\sigma\in{\mathrm{Aut}}_{\Gamma}(L_{V})$.
For instance, one can take for $\Gamma$ the free group $\Gamma={\mathbb{F}}_{2g}=\langle a_{1},\dots,a_{2g}\rangle$, and let $c$ be the product of commutators $[a_{1},a_{2}]\dots[a_{2g-1},a_{2g}]$. In this case, the double $L_{V}$ is the fundamental group of an orientable surface of genus $2g$, and $\Gamma\subset L_{V}$ is the fundamental group of a subsurface of genus $g$ with one boundary component. The variety $V$ itself is the subset of $\Gamma^{2g}$ defined by the equation
$V=\{(x_{1},\dots,x_{2g})|[x_{1},x_{2}]\dots[x_{2g-1},x_{2g}]=[a_{1},a_{2}]\dots[a_{2g-1},a_{2g}]\}.$
The automorphism group ${\mathrm{Aut}}_{\Gamma}(L_{V})$ is isomorphic to the mapping class group of the surface of genus $g$ with one boundary component. This group acts freely on $V=\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ because $\Gamma$-homomorphisms $L_{V}\rightarrow\Gamma$ are injective on $\bar{\Gamma}$. In particular, points of $V$ are in bijection with elements of the mapping class group.
There are very many other test elements in a free group [ST17]. In particular, there exist test elements $c$ for which $\bar{\mathbb{F}}_{n}$ is a rigid vertex of the amalgam ${\mathbb{F}}_{n}*_{c}\bar{\mathbb{F}}_{n}$ (take a test element $c$ whose cyclic reduction contains all subwords of length 3 and apply [CM15]). In this case, ${\mathrm{Aut}}_{\Gamma}(L_{V})$ still acts freely and transitively on $V$, but it is virtually cyclic.
5.3 Grushko decomposition for a homogeneous variety
Proposition 5.6.
If $V$ is homogeneous and irreducible, the Grushko decomposition of $L_{V}$ relative to $\Gamma$ has the form $L_{V}=L^{0}_{V}*{\mathbb{F}}_{r}$ with $\Gamma\subset L^{0}_{V}$, and $L^{0}_{V}$ is the group of functions of a homogeneous irreducible variety.
Proof.
Let $L_{V}=L^{0}_{V}*G_{1}*\dots*G_{n}*{\mathbb{F}}_{r}$ be the Grushko decomposition of $L_{V}$ relative to $\Gamma$, with $\Gamma\subset L^{0}_{V}$, and note that ${\mathrm{Aut}}_{\Gamma}(L_{V})$ permutes the conjugacy classes of the factors $G_{1},\dots,G_{n}$. Choose any $\Gamma$-homomorphism $h_{0}\in\mathrm{Hom}_{\Gamma}(L^{0}_{V},\Gamma)$, and extend it to a $\Gamma$-homomorphism $h\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ that kills $G_{1},\dots,G_{n}$. Since $V$ is homogeneous, the set of $\Gamma$-homomorphisms $\{h\circ\alpha\mid\alpha\in{\mathrm{Aut}}_{\Gamma}(L_{V})\}$ is discriminating by Corollary 5.4. But all of them kill $G_{1}$,…, $G_{n}$. This prevents this family from being discriminating unless $n=0$.
As a subgroup of $L_{V}$, $L_{V}^{0}$ is fully residually $\Gamma$. Thus by Lemma 2.12, the variety associated to $L_{V}^{0}$ is irreducible. To prove that this variety is homogeneous, it is enough by Lemma 5.3 to check that ${\mathrm{Aut}}_{\Gamma}(L_{V}^{0})$ acts transitively on $\mathrm{Hom}_{\Gamma}(L^{0}_{V},\Gamma)$. Given $h_{0},h^{\prime}_{0}\in\mathrm{Hom}_{\Gamma}(L^{0}_{V},\Gamma)$, extend them to $h,h^{\prime}\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ by killing the factor ${\mathbb{F}}_{r}$. By homogeneity of $V$, there exists $\alpha\in{\mathrm{Aut}}_{\Gamma}(L_{V})$ such that $h^{\prime}=h\circ\alpha$. Since $\alpha$ preserves $L^{0}_{V}$, we get that $h^{\prime}_{0}=h_{0}\circ\alpha_{|L_{V}^{0}}$. ∎
5.4 JSJ decomposition for a homogeneous variety
Proposition 5.7.
Let $V$ be an irreducible variety and let $\Lambda$ be a standard JSJ decomposition of its group of functions $L_{V}$.
If $V$ is homogeneous, then all the edge groups of $\Lambda$ are cyclic and every $\Gamma$-homomorphism $h\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ is injective in restriction to all edge groups and rigid groups of $\Lambda$.
Proof.
By Proposition 5.6, it is enough to show the result in the case where $L_{V}$ is one ended.
Let ${\mathrm{Aut}}^{0}_{\Gamma}(L_{V})$ be the subgroup of finite index of ${\mathrm{Aut}}_{\Gamma}(L_{V})$ consisting of automorphisms that act as the identity on the graph $\Lambda$ (recalling that the canonical JSJ tree is invariant under automorphisms).
Let $h\in{\mathrm{Hom}}_{\Gamma}(L_{V},\Gamma)$. By Corollary 5.4 and Lemma 5.5, the family $\{h\circ\alpha\mid\alpha\in{\mathrm{Aut}}^{0}_{\Gamma}(L_{V})\}$ is discriminating. Thus, by Remark 2.10, there exists a sequence of automorphisms $\alpha_{i}\in{\mathrm{Aut}}^{0}_{\Gamma}(L_{V})$ such that $h\circ\alpha_{i}$ is a discriminating sequence.
Let $R\subset L_{V}$ be a rigid vertex group of $\Lambda$, and $E_{1},\dots,E_{p}\subset R$ the collection of incident edge groups. Since each $E_{k}$ is abelian, $h\circ\alpha_{i}(E_{k})$ is cyclic. Up to removing finitely many terms in the sequence we may assume that $h\circ\alpha_{i}(E_{k})$ is non-trivial for all $i,k$. Since for all $i$, $\alpha_{i}(E_{k})$ is conjugate to $E_{k}$ in $L_{V}$, there exist maximal cyclic groups $\langle u_{k}\rangle\subset\Gamma$ such that for all $i$, $h\circ\alpha_{i}(E_{k})$ is conjugate in $\langle u_{k}\rangle$.
We now view $\Gamma$ as hyperbolic relative to $\langle u_{1}\rangle,\dots,\langle u_{k}\rangle$, and we look at an action of $\Gamma$ on a space with a corresponding collection of disjoint horoballs as in [Bow12]. Assume first that the homomorphisms $h\circ\alpha_{i}$ are pairwise distinct modulo conjugation in $\Gamma$. Since $R$ is non-abelian, $h\circ\alpha_{i}(R)$ is non-parabolic for $i$ large enough. By [BS08, Theorem 1.2] (whose proof also works for a discriminating sequence of homomorphisms), we get a non-trivial action of $R$ on an ${\mathbb{R}}$-tree relative to $E_{1},\dots,E_{p}$ with abelian arc stabilizers. Since abelian subgroups of $L_{V}$ are finitely generated, the hypotheses of [Gui08, Theorem 5.1, Corollary 5.2] are satisfied, and this action yields an abelian splitting of $R$ relative to $E_{1},\dots,E_{p}$. Taking a tree of cylinders for commutation as in [GL11, Proposition 6.3], one gets a non-trivial abelian splitting of $R$ relative to $E_{1},\dots,E_{p}$ and to all non-cyclic abelian subgroups of $R$, contradicting its rigidity.
Thus, up to conjugating and taking a subsequence, we may assume that $(h\circ\alpha_{i})_{|R}$ does not depend on $i$. Since it is discriminating, this implies that $(h\circ\alpha_{i})_{|R}$ is injective, hence so is $h_{|R}$.
This shows that $h$ is injective on rigid vertex groups of $\Lambda$. It follows that every edge group $E$ of $\Lambda$ is cyclic because every edge of $\Lambda$ is adjacent to a rigid or surface vertex. Since $h\circ\alpha_{i}$ is discriminating and $\Gamma$ is torsion-free, $h\circ\alpha_{i}$ is injective on $E$ for $i$ large enough. Since $\alpha_{i}$ preserves the conjugacy class of $E$, this shows that $h$ is injective on $E$. ∎
Lemma 5.8.
Let $V$ be a homogeneous irreducible variety over $\Gamma$. Then a standard JSJ decomposition $\Lambda$ of $L_{V}$ relative to $\Gamma$ satisfies the following:
  1. 1.
    in each abelian vertex group, the subgroup generated by incident edge groups is cyclic;
  2. 2.
    $G_{v_{\Gamma}}=\Gamma$;
  3. 3.
    for every non-trivial edge group $G_{e}$ incident on $v_{\Gamma}$, $G_{e}$ is maximal cyclic in $L_{V}$
  4. 4.
    If $e,e^{\prime}$ are distinct edges incident on $v_{\Gamma}$, then $G_{e}$ does not commute with any $L_{V}$-conjugate of $G_{e^{\prime}}$. In particular, for every abelian vertex $v$ adjacent to $v_{\Gamma}$, there is at most one edge joining $v$ to $v_{\Gamma}$.
Proof.
Using Proposition 5.6, we may assume without loss of generality that $L_{V}$ is freely indecomposable relative to $\Gamma$, and that $L_{V}=\pi_{1}(\Lambda)$ is the canonical JSJ decomposition of $L_{V}$.
Let ${\mathrm{Aut}}^{0}_{\Gamma}(L_{V})$ be the subgroup of ${\mathrm{Aut}}_{\Gamma}(L_{V})$ consisting of automorphisms that act as the identity on the graph $\Lambda$ and whose restriction to each edge group is a conjugation. The edge groups of $\Lambda$ are cyclic by Proposition 5.7 so this is a finite index subgroup of ${\mathrm{Aut}}_{\Gamma}(L_{V})$.
Let $h\in{\mathrm{Hom}}_{\Gamma}(L_{V},\Gamma)$. By Corollary 5.4 and Lemma 5.5, there exists a sequence of automorphisms $\alpha_{i}\in{\mathrm{Aut}}^{0}_{\Gamma}(L_{V})$ such that $h\circ\alpha_{i}$ is a discriminating sequence.
To prove (1), let $A$ be an abelian vertex group of $\Lambda$, and $B\subset A$ be the subgroup generated by incident edge groups. Each $\alpha_{i}$ preserves the conjugacy class of $A$ and acts by conjugation on each incident edge group. Using CSA property, it easily follows that $\alpha_{i}$ acts by conjugation on $B$. Thus, $\ker(h\circ\alpha_{i})_{|B}=\ker h_{|B}$ and does not depend on $i$. Since this is a discriminating sequence, $h_{|B}$ is injective. Since $\Gamma$ is hyperbolic, it follows that $B$ is cyclic.
For (2), recall that $G_{v_{\Gamma}}$ denotes the vertex group that contains $\Gamma$, and consider any $h\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$. The vertex group $G_{v_{\Gamma}}$ is rigid by Proposition 3.3 (3). By Proposition 5.7, $h$ is injective on $G_{v_{\Gamma}}\supset\Gamma$. On the other hand, being a $\Gamma$-morphism, $h$ is the identity on $\Gamma$, so $G_{v_{\Gamma}}=\Gamma$.
To prove (3), consider an edge $e$ incident on $v_{\Gamma}$. By Proposition 5.7, $G_{e}$ is infinite cyclic. By Lemma 3.3(5), $G_{e}$ is maximal abelian hence maximal cyclic in $G_{v_{\Gamma}}=\Gamma$. Assuming that $G_{e}$ is not maximal cyclic in $L_{V}$, consider $c\in L_{V}$ and $k>1$ such that $G_{e}=\langle c^{k}\rangle$. Choose some $h\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ (there exists one because $L_{V}$ is a $\Gamma$-limit group), and view it as a retraction $L_{V}\rightarrow\Gamma$. Since $c^{k}\in\Gamma$, we have $c^{k}=h(c^{k})=h(c)^{k}$ contradicting that that $G_{e}$ is maximal cyclic in $\Gamma$.
To prove (4), consider two distinct edges $e,e^{\prime}$ incident on $v_{\Gamma}$ whose edge groups have commuting conjugates in $L_{V}$. View $G_{e}$ and $G_{e^{\prime}}$ as subgroups of $\Gamma$. Using a retraction $L_{V}\rightarrow\Gamma$ as above, we see that $G_{e}$ and $G_{e^{\prime}}$ have commuting conjugates in $\Gamma$. By Lemma 3.3(6), this implies $e=e^{\prime}$. ∎
6 Extension of centralizers and algebraic subgroups
The goal of this section is Theorem 6.8 showing that the JSJ decomposition associated to a connected algebraic group $V$ has no rigid vertex except for $v_{\Gamma}$. We will first associate to $V$ an algebraic subgroup $W_{0}\subset V$ whose associated limit group is obtained by replacing some abelian vertex groups by the cyclic subgroup generated by incident edge groups. This will be used to slightly simplify some arguments including the description of $L_{V\times V}$.
6.1 An algebraic subgroup associated to an extension of centralizers
Let $V$ be a connected algebraic group. Let $\Lambda$ be a standard JSJ decomposition of $L_{V}$ (we don’t assume that $L_{V}$ is one-ended).
Denote by $A_{1},\dots,A_{q}$ the collection of abelian vertex groups in $\Lambda$. Using Lemma 5.8(1), for each $i\leq q$, denote by $\langle c_{i}\rangle$ the smallest direct factor of $A_{i}$ containing all incident edge groups. Let $L_{W}\subset L_{V}$ be the subgroup obtained by replacing each abelian vertex group $A_{i}$ by the cyclic group $\langle c_{i}\rangle$.
We will apply the following lemma to the $\Gamma$-homomorphism $h_{e}\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ corresponding to the neutral element $e\in V$.
Lemma 6.1.
Let $V$ be a connected  algebraic group as above.
Then for any $h\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$, each abelian vertex group $A_{i}$ splits as $A_{i}=\langle c_{i}\rangle\oplus\ker h_{|A_{i}}$.
The lemma actually holds in fact for any homogeneous and irreducible variety.
Proof.
Let $h\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ be any $\Gamma$-homomorphism. We know by Proposition 5.7 that for all $i\leq q$, $h$ is injective on $\langle c_{i}\rangle$ and that $h$ has cyclic image. This does not imply the result because it might happen that $h(\langle c_{i}\rangle)\subsetneqq h(A_{i})$.
For each $i\leq q$, choose an arbitrary decomposition $A_{i}=\langle c_{i}\rangle\oplus A^{\prime}_{i}$, and consider the retraction $r:L_{V}\rightarrow L_{W}$ killing $A^{\prime}_{i}$ for all $i$. Now consider the $\Gamma$-homomorphism $h^{\prime}=h\circ r$. Note that for each $i\leq q$, $\ker h^{\prime}_{|A_{i}}=A^{\prime}_{i}$ because $h^{\prime}$ is injective on $\langle c_{i}\rangle$ by Proposition 5.7. By homogeneity, there exists $\alpha\in{\mathrm{Aut}}_{\Gamma}(L_{V})$ such that $h=h^{\prime}\circ\alpha$.
Fix an abelian vertex group $A_{i}$. Up to precomposing $h$ by an inner automorphism (which does not change its kernel), we may assume that $\alpha(A_{i})=A_{j}$ for some $j\leq q$. Then $\ker h_{|A_{i}}=\alpha^{-1}(\ker h^{\prime}_{|A_{j}})=\alpha^{-1}(A^{\prime}_{j})$. Since $\alpha^{-1}(\langle c_{j}\rangle)=\langle c_{i}\rangle$, the decomposition $A_{j}=\langle c_{j}\rangle\oplus A^{\prime}_{j}$ gives a decomposition $A_{i}=\langle c_{i}\rangle\oplus(\ker h_{|A_{i}})$. ∎
For each $i\leq q$, denote by $N_{i}=\ker h_{e|A_{i}}$ where $h_{e}\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ is the $\Gamma$-homomorphism corresponding to the neutral element $e\in V$. By Lemma 6.1, $A_{i}=\langle c_{i}\rangle\oplus N_{i}$.
We now define the variety and limit group obtained by trimming a collection of abelian groups in $L_{V}$ (see Figure 2). Given a subset ${\mathcal{A}}\subset\{1,\dots,q\}$, denote by $L_{W_{\mathcal{A}}}\subset L_{V}$ the subgroup obtained by replacing, for each $i\in{\mathcal{A}}$, the abelian vertex group $A_{i}$ in ${\mathcal{A}}$ by the cyclic group $\langle c_{i}\rangle$. Thus $L_{V}$ can be viewed as a multiple extension of centralizers $L_{V}=L_{W_{\mathcal{A}}}*_{\langle c_{i}\rangle}(\langle c_{i}\rangle\oplus N_{i})$ (indexed by $i\in{\mathcal{A}}$). Consider the retraction $r:L_{V}\rightarrow L_{W_{\mathcal{A}}}$ killing $N_{i}$ for all $i\in{\mathcal{A}}$, and denote by $j:L_{W_{\mathcal{A}}}\rightarrow L_{V}$ the inclusion. Viewing the set $\mathrm{Hom}_{\Gamma}(L_{W_{\mathcal{A}}},\Gamma)$ as an algebraic variety $W_{\mathcal{A}}$, the precomposition by $r$ and $j$ respectively yield an embedding $r^{*}:W_{\mathcal{A}}={\mathrm{Hom}}_{\Gamma}(L_{W_{\mathcal{A}}},\Gamma)\rightarrow V=\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ and a retraction $j^{*}:V\rightarrow W_{\mathcal{A}}$ such that $j^{*}\circ r^{*}=(r\circ j)^{*}=\mathrm{id}_{W_{\mathcal{A}}}$.
Figure 2: Trimming ${\mathcal{A}}=\{A_{1},A_{2},A_{3}\}$. Round vertices carry abelian groups, square vertices are rigid. The abelian vertex $\langle c_{1}\rangle$ disappears when we pass to the minimal splitting.
Definition 6.2 (Trimming a collection of abelian groups).
We say that $W_{\mathcal{A}}$ or $L_{W_{\mathcal{A}}}$ are obtained from $V$ or $L_{V}$ by trimming the collection of abelian vertex groups ${\mathcal{A}}$.
The following lemma says that the subset $r^{*}(W)$ is an algebraic subgroup of $V$ if the collection of abelian vertices of $\Lambda$ corresponding to ${\mathcal{A}}$ is invariant under ${\mathrm{Aut}}_{\Gamma}(L_{V})$.
Lemma 6.3.
Consider a connected algebraic group $V$, and $\Lambda$ a standard JSJ decomposition of $L_{V}$. Consider a collection of abelian vertex groups ${\mathcal{A}}=\{A_{i}\}$ of $\Lambda$, such that the corresponding set of vertices of $\Lambda$ is invariant under ${\mathrm{Aut}}_{\Gamma}(L_{V})$.
Then the image of the embedding $r^{*}:W_{\mathcal{A}}\rightarrow V$ is an algebraic subgroup of $V$.
Proof.
Note that $r^{*}(W_{\mathcal{A}})$ is the set of $\Gamma$-homomorphisms $h\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ that kill $N_{i}$ for all $i\in{\mathcal{A}}$. In particular, it is a subvariety of $V$. Let $w_{1},w_{2}\in r^{*}(W_{\mathcal{A}})$ and $h_{w_{1}},h_{w_{2}}\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ be the corresponding $\Gamma$-homomorphisms.
Let us check that $w_{1}\circledast w_{2}$ is in $r^{*}(W_{\mathcal{A}})$, i.e. that $h_{w_{1}\circledast w_{2}}$ kills every $N_{i}$. We have $h_{w_{i}}=h_{e}\circ\rho^{*}_{w_{i}}$ (see Lemma 4.2).
By Corollary 4.3, we get
$h_{w_{1}\circledast w_{2}}=h_{e}\circ\rho_{w_{1}}^{*}\circ\rho_{w_{2}}^{*}.$
We know that $h_{w_{1}}=h_{e}\circ\rho^{*}_{w_{1}}$ kills $N_{i}$ for all $i\in{\mathcal{A}}$, it thus suffices to prove that $\rho^{*}_{w_{2}}$ permutes the conjugacy classes of the groups $N_{i}$. So fix $i\in{\mathcal{A}}$. Since $\rho^{*}_{w_{2}}$ permutes the vertices associated to ${\mathcal{A}}$, consider $A_{i^{\prime}}\in{\mathcal{A}}$ such that $\rho^{*}_{w_{2}}(A_{i})$ is conjugate to $A_{i^{\prime}}$ and in particular, $\rho(N_{i})$ is conjugate in $A_{i^{\prime}}$. Now $h_{w_{2}}=h_{e}\circ\rho^{*}_{w_{2}}$ kills $N_{i}$, hence $\rho^{*}_{w_{2}}$ sends $N_{i}$ into a conjugate of $\ker(h_{e|A_{i^{\prime}}})=N_{i^{\prime}}$. This shows that $\rho_{w_{2}}^{*}$ permutes the groups $\{N_{i}|i\in{\mathcal{A}}\}$ up to conjugation, and that $h_{w_{1}\circledast w_{2}}$ kills every $N_{i}$ for $i\in{\mathcal{A}}$. Thus $h_{w_{1}\circledast w_{2}}\in r^{*}(W_{\mathcal{A}})$ which concludes the proof that $r^{*}(W_{\mathcal{A}})$ is stable under multiplication.
Since $\rho^{*}_{\check{w}}=(\rho^{*}_{w})^{-1}$ permutes the $N_{i}$’s up to conjugation, $h_{\check{w}}=h_{e}\circ(\rho^{*}_{w})^{-1}$ kills every $N_{i}$ which shows that that $r^{*}(W_{\mathcal{A}})$ is stable under taking inverse. ∎
Lemma 6.4.
Let $V$ be a connected algebraic group $V$, and let $\Lambda$ be a standard JSJ decomposition of $L_{V}$. Let ${\mathcal{A}}_{0}$ be the collection of non-cyclic abelian vertex groups of $\Lambda$ adjacent to $v_{\Gamma}$.
Denote by $L_{W_{0}}$ the group obtained from $L_{V}$ by trimming ${\mathcal{A}}_{0}$. Let $\Lambda_{W_{0}}$ be the decomposition of $L_{W_{0}}$ obtained from $\Lambda$ by replacing, for each $i\in{\mathcal{A}}_{0}$, the group $A_{i}$ by its edge summand $\langle c_{i}\rangle$ and passing to a minimal subgraph of groups.
Then $\Lambda_{W_{0}}$ is a standard JSJ decomposition of $L_{W_{0}}$.
Remark 6.5.
For each edge group $G_{e}$ incident on $v_{\Gamma}$, the maximal abelian subgroup of $L_{W_{0}}$ containing $G_{e}$ is the cyclic group $G_{e}$ itself (see Lemma 5.8(3)).
Proof.
We first claim that every abelian splitting $L_{W_{0}}=\pi_{1}(\Delta)$ relative to $\Gamma$ and to non-cyclic abelian groups extends to a splitting $\hat{\Delta}$ of $L_{V}$ relative to $\Gamma$ and to its non-cyclic abelian groups. Indeed, write $L_{V}$ as an extension of centralizers $L_{V}=L_{W_{0}}*_{\langle c_{i}\rangle}A_{i}$. To refine this splitting using $\Delta$, we need to check that $c_{i}$ is elliptic in $\Delta$. This is the case: $\langle c_{i}\rangle$ is conjugate in $\Gamma$ since $v_{i}$ is adjacent to $v_{\Gamma}$. This proves the claim.
Since edge groups of $\Lambda$ are elliptic in $\hat{\Delta}$, this implies that edge groups of $\Lambda_{W_{0}}$ are elliptic in $\Delta$. This shows that $\Lambda_{W_{0}}$ is universally elliptic.
To deduce that $\Lambda_{W_{0}}$ is an abelian JSJ decomposition of $L_{W_{0}}$ relative to $\Gamma$ and to its non-cyclic abelian subgroups, it suffices to check that for each vertex $v$ of $\Lambda_{W_{0}}$, $G_{v}$ has a trivial JSJ decomposition relative to $\Gamma$ and to its incident edge groups by Lemma 3.4. This is clear for abelian vertices, and also for non-abelian vertices having the same incident edges in $\Lambda_{W_{0}}$ and $\Lambda$.
It may happen that some edges of $\Lambda$ are not in $\Lambda_{W_{0}}$ because of the need to pass to the minimal splitting. To make the graph of groups minimal, one removes edges $e$ with a valence $1$ endpoint $v$ such that $G_{e}=G_{v}$, and repeat this operation. In our situation, these edges are exactly the edges of $\Lambda$ joining $v_{\Gamma}$ to a terminal abelian vertex. Once these edges have been deleted, then the graph of groups is minimal because the vertex group at $v_{\Gamma}$ is non-abelian. We conclude that the only non-abelian vertex of $\Lambda_{W_{0}}$ whose peripheral structure may have changed is $v_{\Gamma}$. Since we are looking at splittings relative to $\Gamma$, this concludes the proof that $\Lambda_{W_{0}}$ is a JSJ decomposition.
One easily checks that $\Lambda_{W_{0}}$ is actually a standard JSJ decomposition of $L_{W_{0}}$ by applying Lemma 3.2 to the complement of the set of edges with trivial stabilizer. ∎
6.2 Coordinate group of the double
The goal of this subsection is to show that $L_{W_{0}\times W_{0}}$ isomorphic to $L_{W_{0}}*_{\Gamma}L_{W_{0}}$ (Corollary 6.7). Note that this would not be true if we had not trimmed abelian groups neighbouring $v_{\Gamma}$. In view of Lemma 2.13, we rely on the following lemma.
Lemma 6.6.
Let $L$ be the function group of an irreducible variety $W$. Assume that $L$ has a cyclic splitting $\Theta$ such that
  1. 1.
    there is a vertex $v_{\Gamma}$ with vertex group $\Gamma$
  2. 2.
    there is no edge with non-trivial stabilizer joining $v_{\Gamma}$ to itself
  3. 3.
    $\Theta$ is $1$-acylindrical: every non-trivial element fixes at most one edge.
Then the coordinate group of $W\times W$ is the double $L*_{\Gamma}\bar{L}$ where $\bar{L}$ is an isomorphic copy of $L$.
Proof.
Denote by $D$ the double $L*_{\Gamma}\bar{L}$. By Lemma 2.13, the coordinate group of $W\times W$ is the maximal residually $\Gamma$ quotient of $D$. Thus, it suffices to prove that $D$ is a $\Gamma$-limit group.
For any $g\in L$, denote by $\bar{g}$ its copy in $\bar{L}\subset D$. Let $\varphi:L*_{\Gamma}\bar{L}\rightarrow L$ be the morphism that is the identity on the left factor $L$, and maps $\bar{g}\in\bar{L}$ to $g\in L$.
One can write $D$ as a the fundamental group of the graph of groups $\Theta^{2}$ obtained by gluing $\Theta$ with a copy $\bar{\Theta}$ of itself along the vertex $v_{\Gamma}$. One easily checks that $\Theta^{2}$ is $2$-acylindrical: any segment of length 3 in the Bass-Serre tree of $\Theta^{2}$ has a subsegment of length $2$ in a Bass-Serre tree of $\Theta$ or $\bar{\Theta}$ by Assumption 2, so one concludes using $1$-acylindricity of $\Theta$ and $\bar{\Theta}$.
Non-trivial edge groups of $\Theta$ are maximal abelian in $L$ by $1$-acylindricity of $\Theta$. By [CG05, Corollary A.8], $D$ is CSA.
The morphism $\varphi$ is injective on all vertex groups of $\Theta^{2}$. Since $\Theta^{2}$ is a graph of groups whose non-trivial edge groups are maximal abelian in neighbouring vertex groups, and since $L$ is a $\Gamma$-limit group, it follows that $D$ is a $\Gamma$-limit group by [Sel09, Theorem 1.31], or [CG05, Prop 4.11] which adapts immediately to the context of $\Gamma$-limit groups. ∎
Corollary 6.7.
Let $V$ be a connected algebraic group, and $W$ its ${\mathcal{A}}_{0}$-trimming as in Lemma 6.4.
Then $L_{{W_{0}}\times{W_{0}}}=L_{W_{0}}*_{\Gamma}L_{W_{0}}$.
Proof.
Let $\Lambda_{W_{0}}$ be the decomposition of $L_{W_{0}}$ as in Lemma 6.4. Let $\Theta$ be the graph of groups obtained from $\Lambda_{W_{0}}$ by collapsing all edges that are not adjacent to $v_{\Gamma}$.
We apply Lemma 6.6. By Lemma 5.8 (2) and recalling that $\Lambda_{W_{0}}$ is bipartite, we get that the first two assumptions of the lemma hold. Thus it suffices to check that $\Theta$ is $1$-acylindrical.
To prove this consider $e\neq e^{\prime}$ two edges with common vertex $w$ in the Bass-Serre tree of $\Theta$ and with $G_{e}\cap G_{e^{\prime}}\neq\{1\}$. Then $G_{e}=G_{e^{\prime}}$ by Remark 6.5. The two edges $e$ and $e^{\prime}$ are not in the same $G_{w}$-orbit because being maximal abelian in $L_{W_{0}}$, $G_{e}$ is malnormal in $G_{w}$. Downstairs, the image of $w$ need not be equal to $v_{\Gamma}$ but the images of $e$ and $e^{\prime}$ in $\Theta$ are two distinct edges incident on $v_{\Gamma}$ whose edge groups are conjugate in $L_{W_{0}}$, contradicting Lemma 5.8 (4). This shows that $\Theta$ is $1$-acylindrical and concludes the proof. ∎
6.3 Rigid vertices in the JSJ decomposition
Theorem 6.8.
Let $V$ be a connected algebraic group, and $\Lambda_{V}$ be a standard JSJ decomposition of $L_{V}$.
Then $\Lambda_{V}$ has no rigid vertex other than the vertex $v_{\Gamma}$ with vertex group $\Gamma$.
Proof of Theorem 6.8.
Let $\Lambda$ be a JSJ decomposition of $L_{V}$.
Let $L_{W_{0}}$ and $\Lambda_{W_{0}}$ be obtained by trimming ${\mathcal{A}}_{0}$ as in Lemma 6.4. It suffices to show that $\Lambda_{W_{0}}$ has no rigid vertex distinct from $v_{\Gamma}$.
By Corollary 6.7, $L_{{W_{0}}\times{W_{0}}}$ is the double $D=L_{W_{0}}*_{\Gamma}\bar{L}_{W_{0}}$.
The multiplication law $\mu:W_{0}\times W_{0}\rightarrow W_{0}$ induces a dual morphism $\mu^{*}:L_{W_{0}}\rightarrow D$. Denote by $h_{e}:L_{W_{0}}\rightarrow\Gamma$ the $\Gamma$-homomorphism corresponding to the identity element $e\in{W_{0}}$. The map $\Psi:w\in{W_{0}}\mapsto(e,w)\in{W_{0}}\times{W_{0}}$ induces a dual morphism $\Psi^{*}:D\rightarrow L_{W_{0}}$. It coincides with $h_{e}$ on $L_{W_{0}}$, and maps $\bar{g}\in\bar{L}_{W_{0}}$ to $g$. We will also use the symmetric map $\Psi^{\prime}:w\mapsto(w,e)$.
Since $\mu\circ\Psi=\mathrm{id}_{W_{0}}$, we get $\Psi^{*}\circ\mu^{*}=\mathrm{id}_{L_{W_{0}}}$ so $\mu^{*}$ is injective. Let $\Lambda_{D}$ be the splitting of $D$ obtained by gluing two copies $\Lambda_{W_{0}},\bar{\Lambda}_{W_{0}}$ of $\Lambda_{{W_{0}}}$ along $v_{\Gamma}$.
Fix a rigid vertex group $R$ of $\Lambda_{W_{0}}$. The monomorphism $\mu^{*}$ gives an action of $L_{W_{0}}$ on the Bass-Serre tree of $\Lambda_{D}$. In this action, $R$ is elliptic thus $\mu^{*}(R)$ is conjugate into a vertex group $G_{u}$ of $\Lambda_{D}$. If $u$ corresponds to a vertex coming from $\Lambda_{W_{0}}$ (allowing $u=v_{\Gamma}$), then $\Psi^{*}(\mu^{*}(R))$ is conjugate in $h_{e}(G_{u})\subset\Gamma$. Since $\Psi^{*}\circ\mu^{*}=\mathrm{id}_{L_{V}}$, we get that $R$ is conjugate in $\Gamma$. If $u$ comes from $\bar{\Lambda}_{W_{0}}$ we proceed similarly using $\Psi^{\prime*}$ instead of $\Psi^{*}$. This concludes the proof of the theorem. ∎
7 Bounded stretch and application to vertices of surface type
7.1 Length and stretch
Let $L$ be a $\Gamma$-group with a finite $\Gamma$-generating set $X_{1},\ldots,X_{n}$, i. e. such that $L=\langle\Gamma,X_{1},\dots,X_{n}\rangle$.
Definition 7.1.
Given $f\in L$, we define the length $\ell(f)$ of $f$ to be the minimal length of a word on the alphabet $\{X^{\pm 1}_{1},\ldots,X^{\pm 1}_{n}\}\cup(\Gamma\setminus\{1\})$ representing $f$.
Remark 7.2.
The length of $f$ should be thought of as an analogue of the degree of a polynomial (although the length of $aX_{1}X_{2}bX_{2}^{-2}$ is 6 while its degree should rather be 4).
Definition 7.3.
Let $\alpha\in{\mathrm{Aut}}_{\Gamma}(L)$. We define the stretch of $\alpha$ (with respect to the $\Gamma$-generating set $X_{1},\dots,X_{n}$ of $L$) as
${\mathrm{Lip}}(\alpha)=\sup_{f\in L\setminus\{1\}}\frac{\ell(\alpha(f))}{\ell(f)}$
$=\max\{1,\ell(\alpha(X_{1})),\dots,\ell(\alpha(X_{n}))\}$
Remark 7.4.
The equivalence between the two definitions follows from the fact that $\Gamma\cup\{X_{1},\dots,X_{n}\}$ is a generating set of $L$, so the stretch of $\alpha$ is bounded by the length of their images.
The following lemma is a key observation. It generalizes the classical fact of algebraic geometry saying that if ${\mathbb{G}}\curvearrowright X$ is an algebraic action of an algebraic group on an algebraic variety $X$, for every function $f$ on $X$, the ${\mathbb{G}}$-orbit of $f$ spans a finite dimensional vector space ([Bri09]). If $X$ is the affine space, this means that the ${\mathbb{G}}$-orbit of a polynomial $f$ has bounded degree.
We say that an algebraic group $G$ acts algebraically on a variety $X$ if the map $G\times X\rightarrow X$ is an algebraic map. Given $g\in G$, we denote by $\tau_{g}:x\in X\mapsto g.x\in X$, and by $\tau_{g}^{*}\in{\mathrm{Aut}}_{\Gamma}(L_{X})$ the automorphism sending $f\in L_{X}$ to $f\circ\tau_{g}:z\mapsto f(g.x)$.
Lemma 7.5 (Bounded stretch).
Let $G$ be an algebraic group acting algebraically on an algebraic variety $X$.
Then there exists a constant $C>0$ such that for all $g\in G$, ${\mathrm{Lip}}(\tau_{g}^{*})\leq C$.
Equivalently, for every function $f\in L_{X}$, and every $g\in G$, $\ell(f\circ\tau_{g})\leq C\ell(f)$.
Proof.
View $G$ (resp $X$) as a subset of $\Gamma^{n}$ (resp $\Gamma^{p}$), and denote by $Z_{1},\dots,Z_{n}:G\rightarrow\Gamma$ (resp $X_{1},\dots,X_{p}:X\rightarrow\Gamma$) the coordinate functions – sending $(z_{1},\dots,z_{n})\in G$ to $z_{i}$. The action map $\alpha:G\times X\rightarrow X$ can be written as $(\alpha_{1},\dots,\alpha_{p})$ where each $\alpha_{i}$ is a word on $n+p$ variables $\alpha_{i}(z_{1},\dots,z_{n},x_{1},\dots,x_{p})$.
Given $g\in G$, the map $\tau_{g}:X\rightarrow X$ is defined by $\tau_{g}(.)=\alpha(g,.)$. Thus, $\tau_{g}^{*}(X_{i})=X_{i}\circ\tau_{g}=X_{i}(\alpha(g,.))=\alpha_{i}(g,.)=\alpha_{i}(g_{1},\dots,g_{n},X_{1},\dots,X_{p})\in L_{X}$. Thus $\ell(\tau_{g}^{*}(X_{i}))$ is bounded by the length of $\alpha_{i}$ as a word in $n+p$ variables. This bound does not depend on $g$, which proves the lemma. ∎
Given an element $g$ acting isometrically on a metric space $Y$, we denote by $||g||_{Y}=\inf_{y\in Y}d(y,gy)$ its translation length. By the triangle inequality, we obtain:
Lemma 7.6.
Let $G$ be an algebraic group acting algebraically on an algebraic variety $X$. Consider an action of $L_{X}$ by isometries on a metric space $Y$, such that $\Gamma$ has bounded orbits.
Then there exists a constant $K$ such that for all $f\in L_{X}$ and $g\in G$, $||\tau_{g}^{*}(f)||_{Y}\leq K\ell(f)$.
Proof.
Endow $L_{X}=\langle\Gamma,X_{1},\dots,X_{p}\rangle$ with the word metric $\ell$ on the infinite generating set $\{X_{1},\dots,X_{p}\}\cup\Gamma\setminus\{1\}$. Fix a basepoint $*\in Y$. The orbit map $L_{X}\to Y$ given by $f\mapsto f.*$ is $M$-Lipschitz where $M$ is an upper bound on the diameter of $\Gamma.*$ and on the distances $d_{Y}(*,X_{i}.*)$. Thus, using Lemma 7.5, for all $f\in L_{V}$ and $g\in G$,
$d_{Y}(*,f\circ\tau_{g}.*)\leq M\ell(f\circ\tau_{g})\leq MC\ell(f).$
If $V$ is an algebraic group, we can apply the above result with $G=V$ acting on $X=V$ by right multiplication, i.e. $G\times X\to X$ is given by $(g,x)\mapsto x\circledast g$. In this case $\tau_{g}=\rho_{g}$ so we get
Corollary 7.7.
Let $V$ be an algebraic group. Consider an action of $L_{V}$ by isometries on a metric space $Y$, such that $\Gamma$ has bounded orbits.
Then there exists a constant $K$ such that for all $f\in L_{V}$, and all $v\in V$, $||\rho_{v}^{*}(f)||_{Y}\leq K\ell(f)$.
We will also use Lemma 7.6 for the action of the algebraic group $G=V\times V$ on the variety $V$ by left and right multiplication given by $(a,b)\cdot x=a\circledast x\circledast b$ (formally, this is rather an action of $V\times\check{V}$ on $V$ where $\check{V}$ is the algebraic group $V$ with the multiplication law $\check{\mu}$ defined by $\check{\mu}(\underline{v}_{1},\underline{v}_{2})=\mu(\underline{v}_{2},\underline{v}_{1})$).
For $a,b\in V$ we denote by $\Delta_{a,b}:V\to V;v\mapsto a\circledast v\circledast b$. In this setting Lemma 7.6 gives:
Corollary 7.8.
Let $V$ be an algebraic group. Consider an action of $L_{V}$ by isometries on a metric space $Y$, such that $\Gamma$ has bounded orbits.
Then for all $f\in L_{V}$, there exists a constant $C_{f}$ such that for all $a,b\in V$, $||\Delta_{a,b}^{*}(f)||_{Y}\leq C_{f}$. ∎
7.2 Surfaces in the JSJ decomposition
The goal of this section is the following result.
Theorem 7.9.
Let $V$ be a connected algebraic group with coordinate group $L_{V}$, and $\Lambda$ be a standard JSJ decomposition of $L_{V}$.
Then $\Lambda$ has no vertex of surface type whose corresponding surface has infinite mapping class group.
Proof.
Let $Q$ be a vertex group of surface type of $\Lambda$. Let $V_{0}\subset V$ be the finite index subgroup consisting of elements $v$ such that $\rho_{v}^{*}$ preserves the conjugacy class of $Q$ and of incident edge groups (we don’t claim that $V_{0}$ is an algebraic subgroup of $V$).
Given $v\in V_{0}$, consider $g_{v}\in L_{V}$ such that ${\mathrm{ad}}_{g_{v}}\circ\rho_{v}^{*}(Q)\subset Q$, and consider the restriction $\sigma_{v}=({\mathrm{ad}}_{g_{v}}\circ\rho_{v}^{*})_{|Q}$. Noting that the class of $\sigma_{v}$ in ${\mathrm{Out}}(Q)$ does not depend on the choice of $g_{v}$, the map $v\mapsto\sigma_{v}$ is a well defined morphism $\sigma:V_{0}\rightarrow{\mathrm{Out}}(Q)$. We view $\sigma_{v}$ as an element of the mapping class group of the underlying surface $\Sigma$.
We claim that this morphism has finite image. Otherwise, since ${\mathrm{Out}}(Q)$ is virtually torsion-free, there exists $v\in V$ such that $\sigma_{v}$ has infinite order. Consider closed curves $\beta,\beta^{\prime}$ in the surface $\Sigma$ underlying $Q$ such that $\beta$ is simple, and the intersection number $i(\beta,\sigma_{v}^{n}(\beta^{\prime}))$ is unbounded.
Let $T_{\beta}$ be the refinement of the JSJ decomposition of $L_{V}$ given by the curve $\beta$ (if $\beta$ is one-sided, this is the splitting dual to the boundary of a Möbius band with core curve $\beta$). Let $g\in Q$ be an element representing $\beta^{\prime}$. Then $||\sigma_{v}^{n}(g)||_{T_{\beta}}=i(\beta,\sigma_{v}^{n}(\beta^{\prime}))$ is unbounded. Since $\sigma_{v}^{n}(g)$ equals $\rho_{v}^{*n}(g)$ up to conjugacy, and $\rho_{v}^{*n}=\rho^{*}_{v^{\circledast n}}$, this contradicts Corollary 7.7 for the space $Y=T_{\beta}$. This concludes the proof that $\sigma$ has finite image.
Let $h\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ be any $\Gamma$-homomorphism. We claim that $h_{|Q}$ is injective. Indeed, let $V_{1}\subset V_{0}$ be the finite index subgroup of $V$ consisting of elements $v\in V_{0}$ such that $\sigma_{v}=\mathrm{id}_{Q}$ in ${\mathrm{Out}}(Q)$. By Lemma 5.5, the family $\{h\circ\rho_{v}^{*}|v\in V_{1}\}$ is discriminating. Since all these morphisms differ by a conjugation on $Q$, $h_{|Q}$ is injective.
Now assume by contradiction that $Q$ has infinite outer automorphism group. This implies that it contains a Dehn twist of infinite order. Extend this Dehn twist to an automorphism $\alpha\in Aut_{\Gamma}(L_{V})$ preserving $Q$, and acting as the identity on a non-abelian subgroup $Q_{1}\subset Q$ (corresponding to the fundamental group of a connected component of the complement of the twisting curve). By homogeneity, and since $V_{1}$ has finite index in $V$, the $h\circ\alpha^{k}$ fall into finitely many $V_{1}$-orbits. Thus we can find distinct indices $k,l\in{\mathbb{N}}$ such that $h\circ\alpha^{k}$ and $h\circ\alpha^{l}$ are in the same $V_{1}$-orbit, i. e. there exists $v\in V_{1}$ such that $h\circ\alpha^{k}=h\circ\alpha^{l}\circ\rho^{*}_{v}$.
Since $\rho^{*}_{v}$ restricts to a conjugation on $Q$, there exists $c\in\Gamma$ such that
$h_{|Q}\circ(\alpha^{k})_{|Q}={\mathrm{ad}}_{c}\circ h_{|Q}\circ(\alpha^{l})_{|Q}$
Since $\alpha_{|Q_{1}}=\mathrm{id}$, $c$ centralizes $h(Q_{1})$, and $c=1$ since $\Gamma$ is torsion-free hyperbolic. Since $h$ is injective on $Q$, $\alpha^{k}_{|Q}=\alpha^{l}_{|Q}$, a contradiction. ∎
8 Classification of the underlying variety
Recall that $h_{e}\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ is the $\Gamma$-homomorphism associated to the neutral element of the algebraic group $V$.
Theorem 8.1.
Let $\Gamma$ be a torsion-free hyperbolic group. Let $V$ be a connected algebraic group over $\Gamma$ with coordinate group $L_{V}$.
Then there exist non-conjugate maximal cyclic groups $\langle c_{1}\rangle,\ldots,\langle c_{l}\rangle$ of $\Gamma$ and $n_{1},\ldots,n_{l}\in{\mathbb{N}}\setminus\{0\}$ such that the algebraic variety $V$ is isomorphic to $\Gamma^{r}\times\langle c_{1}\rangle^{n_{1}}\times\dots\times\langle c_{l}\rangle^{n_{l}}$.
Equivalently, $L_{V}$ is the free product of a free group with a multiple extension of centralizers $L_{W}$ of $\Gamma$ along the $c_{j}$’s as on Figure 3, where the abelian vertex groups are given by $A_{j}=\langle c_{j}\rangle\oplus N_{j}$, with $N_{j}=\ker h_{e|A_{j}}\simeq{\mathbb{Z}}^{n_{j}}$.
Figure 3: The limit group $L_{V}=L_{W}*{\mathbb{F}}_{r}$ of a connected algebraic group.
Proof.
Let $\Lambda$ be a standard JSJ decomposition of $L_{V}$.
By Proposition 5.6, the Grushko decomposition of $L_{V}$ relative to $\Gamma$ has the form $L_{V}=L_{V}^{0}*{\mathbb{F}}_{r}$ with $\Gamma\subset L_{V}^{0}$. Thus, the subgraph $\Lambda_{0}\subset\Lambda$ spanned by vertices and edges with non-trivial vertex or edge group is connected, with $\pi_{1}(\Lambda_{0})=L_{V}^{0}$. By Lemma 5.8, the rigid group of $\Lambda_{0}$ containing $\Gamma$ is equal to $\Gamma$. We denote by $v_{\Gamma}$ the corresponding vertex of $\Lambda_{0}$.
By Theorem 7.9, $\Lambda_{0}$ has no vertex of surface type, and by Theorem 6.8 it has no rigid vertex except $v_{\Gamma}$. Thus, $v_{\Gamma}$ is the only vertex of $\Lambda_{0}$ with non-abelian stabilizer. Moreover, since the Bass-Serre tree of $\Lambda_{0}$ is a tree of cylinders, it is bipartite between abelian and non-abelian vertices. It follows that every abelian vertex $v$ of $\Lambda_{0}$ is adjacent to $v_{\Gamma}$, and by Assertion 4 of Lemma 5.8, there is exactly one edge in $\Lambda_{0}$ joining $v$ to $v_{\Gamma}$. Thus, $\Lambda_{0}$ is a star with center $v_{\Gamma}$. The incident edge groups $\langle c_{j}\rangle\subset\Gamma$ are non-conjugate in $\Gamma$ by ayclindricity at $v_{\Gamma}$ ($\Lambda_{0}$ is a tree of cylinders).
By Assertion 3 of Lemma 5.8, all edge groups of $\Lambda_{0}$ are maximal cyclic in $L_{V}$. This shows that $L_{V}^{0}$ is a multiple extension of centralizers and Lemma 6.1 concludes the proof. ∎
9 Classification of the multiplication law
Theorem 8.1 gives a full description of the underlying variety of an algebraic group $(V,\mu)$ over $\Gamma$. The aim of this section is to complete the classification of algebraic groups by classifying the possible multiplication laws on $V$. Our goal is to find a system of coordinates in which $\mu$ is the standard multiplication law.
Starting with $T_{1},\ldots,T_{n}\in L_{V}$ such that $L_{V}=\langle\Gamma,T_{1},\ldots,T_{n}\rangle$, we get an embedding $V=\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ into $\Gamma^{n}$ via $h\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)\mapsto(h(T_{1}),\dots,h(T_{n}))$. This way, $T_{i}$ becomes the $i$-th coordinate function on $V$, and we call $(T_{1},\dots,T_{n})$ a system of coordinates on $V$.
Such a system induces a system of coordinates $(X_{1},\dots,X_{n},Y_{1},\dots,Y_{n})$ for $V\times V$ as follows. Denoting by $p_{1},p_{2}:V\times V\rightarrow V$ the two projections, we define $X_{i}=T_{i}\circ p_{1}$ and $Y_{i}=T_{i}\circ p_{2}$. Equivalently, $X_{i}=p_{1}^{*}(T_{i})$ and $Y_{i}=p_{2}^{*}(T_{i})$. Viewing $L_{V\times V}$ as a quotient of $L_{V}*_{\Gamma}L_{V}$, $X_{i}$ (resp. $Y_{i}$) is the image of $T_{i}$ under the left (resp. right) embedding of $L_{V}$ into $L_{V\times V}$.
With these coordinates, if we write the multiplication law $\mu:V\times V\rightarrow V$ by
$\begin{bmatrix}x_{1}\\ \vdots\\ x_{n}\end{bmatrix}\circledast\begin{bmatrix}y_{1}\\ \vdots\\ y_{n}\end{bmatrix}=\begin{bmatrix}\mu_{1}(\underline{x},\underline{y})\\ \vdots\\ \mu_{n}(\underline{x},\underline{y})\end{bmatrix},$
then the comultiplication law $\mu^{*}:L_{V}\rightarrow L_{V\times V}$ is defined on the generating set by
$\mu^{*}:\left\{\begin{matrix}T_{1}\mapsto\mu^{*}(T_{1})=\mu_{1}(X_{1},\dots,X_{n},Y_{1},\dots,Y_{n})\\ \vdots\\ T_{n}\mapsto\mu^{*}(T_{n})=\mu_{n}(X_{1},\dots,X_{n},Y_{1},\dots,Y_{n})\end{matrix}\right.$
Our goal is to find coordinates $T_{1},\ldots,T_{n}\in L_{V}$ in which $\mu$ is the standard multiplication law i.e., $\mu^{*}(T_{i})=X_{i}Y_{i}$ for all $i\leq n$.
Remark 9.1.
If we change the coordinate system $(T_{1},\dots,T_{n})$ on $V$, we need to change the coordinate system on $V^{2}$ accordingly. For instance, if we change $T_{i}$ to $T^{\prime}_{i}=T_{i}^{-1}$, then $X_{i}$ changes to $X^{\prime}_{i}=p_{1}^{*}(T^{\prime}_{i})=X_{i}^{-1}$ and similarly, $Y_{i}$ changes to $Y^{\prime}_{i}=Y_{i}^{-1}$. If we change $T_{i}$ to $T^{\prime}_{i}=uT_{i}v$ for some $u,v\in L_{V}$, then $X_{i}$ changes to $X^{\prime}_{i}=u^{(X)}X_{i}v^{(X)}$ where $u^{(X)}=p_{1}^{*}(u)\in L_{V^{2}}$ and $v^{(X)}=p_{1}^{*}(v)$ are the copies of $u$ and $v$ under the left embedding of $L_{V}$ into $L_{V^{2}}$; there are similar changes for $Y_{i}$.
Just as for $V\times V$, the system of coordinates $(T_{1},\ldots,T_{n})$ on $V$ also induces a system of coordinates on $V\times V\times V$: we define $X_{i},Z_{i},Y_{i}\in L_{V^{3}}$ (in this order) as the image of $T_{i}$ under the left, middle, and right embedding of $L_{V}$ in $L_{V^{3}}$ respectively. This corresponds to coordinates associated to the embedding $V\times V\times V\subset\Gamma^{n}\times\Gamma^{n}\times\Gamma^{n}$.
Let $(\underline{x},\underline{z},\underline{y})$ be a point of $V\times V\times V$. Associativity then reads
$(\underline{x}\circledast\underline{z})\circledast\underline{y}=\underline{x}\circledast(\underline{z}\circledast\underline{y})\quad\text{i.e.}\quad\mu(\mu(\underline{x},\underline{z}),\underline{y})=\mu(\underline{x},\mu(\underline{z},\underline{y}))$
[…]
where $\mu_{X,Z}\times\mathrm{id}_{Y}:V^{3}\rightarrow V^{2}$ is the map sending $(\underline{x},\underline{z},\underline{y})$ to $(\mu(\underline{x},\underline{z}),\underline{y})$ and $\mathrm{id}_{X}\times\mu_{Z,Y}:V^{3}\rightarrow V^{2}$ maps $(\underline{x},\underline{z},\underline{y})$ to $(\underline{x},\mu(\underline{z},\underline{y}))$.
At the level of function groups, the maps induced by $\mu_{X,Z}\times\mathrm{id}_{Y}$ and $\mathrm{id}_{X}\times\mu_{Z,Y}$ are
$\mu^{*}_{X,Z}\otimes\mathrm{id}^{*}_{Y}:\left\{\begin{array}[]{rcl}L_{V\times V}&\rightarrow&L_{V\times V\times V}\\ X_{i}&\mapsto&\mu_{i}(\underline{X},\underline{Z})\\ Y_{i}&\mapsto&Y_{i}\\ \end{array}\right.,\quad\text{and }\mathrm{id}^{*}_{X}\otimes\mu^{*}_{Z,Y}:\left\{\begin{array}[]{rcl}L_{V\times V}&\rightarrow&L_{V\times V\times V}\\ X_{i}&\mapsto&X_{i}\\ Y_{i}&\mapsto&\mu_{i}(\underline{Z},\underline{Y})\\ \end{array}\right.$
and associativity translates into
$(\mu^{*}_{X,Z}\otimes\mathrm{id}^{*}_{Y})\circ\mu^{*}=(\mathrm{id}^{*}_{X}\otimes\mu^{*}_{Z,Y})\circ\mu^{*}.$
9.1 Coordinates on the groups $L_{V}$, $L_{V^{2}}$ and $L_{V^{3}}$
We have seen in Theorem 8.1 that the underlying variety $V$ of an algebraic group is isomorphic to $\Gamma^{r}\times Z(c_{1})^{n_{1}}\times\dots\times Z(c_{l})^{n_{l}}$. In other words, the group $L_{V}$ can be written as the fundamental group of a graph of groups $\Lambda$ as in Figure 3, i. e. it is the free product of a free group ${\mathbb{F}}_{r}$ with a multiple extension of centralizers $\Gamma*_{\langle c_{j}\rangle}A_{j}$ with $A_{j}=\langle c_{j}\rangle\oplus N_{j}$ and $N_{j}=\ker h_{e|A_{j}}\simeq{\mathbb{Z}}^{n_{j}}$, where $h_{e}\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ is the $\Gamma$-homomorphism corresponding to the neutral element $e\in V$.
We denote by $T_{1},\dots,T_{r}$ a free basis of the free group ${\mathbb{F}}_{r}$, and for all $j\leq l$, we denote by $T_{j,1},\dots,T_{j,n_{j}}$ a basis of the free abelian group $N_{j}$, so that in particular $h_{e}(T_{j,k})=1$ for all $j,k$ (see Figure 4). We say that the $T_{i}$ ($i\leq r$) are the free generators, and the $T_{j,k}$ ($j\leq l$, $k\leq n_{j}$) are the abelian generators.
Let $p_{1},p_{2}:V\times V\rightarrow V$ the two projections. As described above, from the coordinate system $T_{i},T_{j,k}$ on $V$, we get a coordinate system $X_{i}$, $X_{j,k}$, $Y_{i}$, $Y_{j,k}$ on $V\times V$ defined by $X_{i}=p_{1}^{*}(T_{i})$, $X_{j,k}=p_{1}^{*}(T_{j,k})$, $Y_{i}=p_{2}^{*}(T_{i})$ and $Y_{j,k}=p_{2}^{*}(T_{j,k})$ (see Figure 4). Similarly, we use the coordinate system $X_{i}$, $X_{j,k}$, $Z_{i}$, $Z_{j,k}$, $Y_{i}$, $Y_{j,k}$ on $V^{3}$ associated to the three projections $V^{3}\rightarrow V$.
Lemma 9.2.
The function groups $L_{V^{2}}$ of $V\times V$ and $L_{V^{3}}$ of $V\times V\times V$ are both free products of multiple extensions of centralizers $\langle c_{1}\rangle,\ldots,\langle c_{l}\rangle$ of $\Gamma$ with a free group, of rank $2r$ and $3r$ respectively (see Figure 4), where for $j\leq l$, the abelian vertex group of $L_{V^{2}}$ containing $\langle c_{j}\rangle$ is
$A_{j}^{V^{2}}=\langle c_{j}\rangle\oplus\langle X_{j,1},\dots,X_{j,n_{j}}\rangle\oplus\langle Y_{j,1},\dots,Y_{j,n_{j}}\rangle$
$=\langle c_{j}\rangle\oplus N_{j}^{(X)}\oplus N_{j}^{(Y)}\simeq\langle c_{j}\rangle\oplus{\mathbb{Z}}^{2n_{j}}$
and the abelian vertex group of $L_{V^{3}}$ containing $\langle c_{j}\rangle$ is
$A_{j}^{V^{3}}=\langle c_{j}\rangle\oplus\langle X_{j,1},\dots,X_{j,n_{j}}\rangle\oplus\langle Z_{j,1},\dots,Z_{j,n_{j}}\rangle\oplus\langle Y_{j,1},\dots,Y_{j,n_{j}}\rangle$
$=\langle c_{j}\rangle\oplus N_{j}^{(X)}\oplus N_{j}^{(Z)}\oplus N_{j}^{(Y)}\simeq\langle c_{j}\rangle\oplus{\mathbb{Z}}^{3n_{j}}.$
Their respective JSJ decompositions are shown on Figure 4.
Figure 4: The groups $L_{V}$, $L_{V^{2}}$, and $L_{V^{3}}$ with their generating sets, and the corresponding graphs of groups $\Lambda_{V}$, $\Lambda_{V^{2}}$ and $\Lambda_{V^{3}}$.
Proof.
By Lemma 2.13, $L_{V^{2}}$ is the maximal residually $\Gamma$ quotient of the double $D=L_{V}*_{\Gamma}\bar{L}_{V}$. For any $\Gamma$-homomorphism $h\in\mathrm{Hom}_{\Gamma}(D,\Gamma)$, and any $j\leq l$, $h(N_{j})$ and $h(\bar{N}_{j})$ commute with $h(c_{j})=h(\bar{c}_{j})\neq 1$, so commute together. These commutation relations therefore hold in $L_{V^{2}}$. Now the quotient of $D$ by these commutation relations is the fundamental group of the graph of groups on Figure 4. Since the fundamental group of this graph of groups is a $\Gamma$-limit group, it is in particular residually $\Gamma$ and is thus equal to $L_{V^{2}}$. The case of $L_{V^{3}}$ is similar. ∎
9.2 Multiplication law on the abelian part
Recall that we defined a coordinate system $T_{i},T_{j,k}\in L_{V}$ for which, for all $j\leq l$,
$\ker h_{e|A_{j}}=N_{j}=\langle T_{j,1},\ldots,T_{j,n_{j}}\rangle.$
We now show that in these coordinates the abelian part of the multiplication law is standard.
Proposition 9.3.
With the bases chosen above, for all $j\leq l$, we have
$\mu^{*}(T_{j,k})=X_{j,k}Y_{j,k}\in L_{V^{2}}.$
Proof.
Fix $j\leq l$. Since $T_{j,k}$ commutes with $c_{j}$ in $L_{V}$, $\mu^{*}(T_{j,k})$ commutes with $c_{j}$ in $L_{V^{2}}$, and we can write $\mu^{*}(T_{j,k})=c_{j}^{q_{j,k}}u_{j,k}v_{j,k}$ where $u_{j,k}\in N_{j}^{(X)}$ and $v_{j,k}\in N_{j}^{(Y)}$.
Denote by $i_{e},i_{e}^{\prime}:V\to V^{2}$ the maps defined by $v\mapsto(e,v)$ and $v\mapsto(v,e)$. At the level of function groups, $i_{e}^{*}:L_{V^{2}}\rightarrow L_{V}$ is induced by the map $L_{V}*_{\Gamma}L_{V}\rightarrow L_{V}$ which is $h_{e}$ on the left factor and the identity on the right factor (sending $Y_{i}$ and $Y_{j,k}$ to $T_{i}$ and $T_{j,k}$). In particular, $N_{j}^{(X)}\subset\ker i_{e}^{*}$. Similarly, $i_{e}^{\prime*}(X_{i})=T_{i}$, $i_{e}^{\prime*}(X_{j,k})=T_{j,k}$, and $i_{e}^{\prime*}$ coincides with $h_{e}$ on the right factor, so $N_{j}^{(Y)}\subset\ker i_{e}^{\prime*}$.
Since $\mu\circ i_{e}=\mu\circ i^{\prime}_{e}=\mathrm{id}_{V}$, we get that $i_{e}^{*}\circ\mu^{*}=i_{e}^{\prime*}\circ\mu^{*}=\mathrm{id}_{L_{V}}$. Using that $u_{j,k}\in N_{j}^{(X)}$ lies in $\ker i^{*}_{e}$ we get
$T_{j,k}=i_{e}^{*}(\mu^{*}(T_{j,k}))=i_{e}^{*}(c_{j}^{q_{j,k}}u_{j,k}v_{j,k})=c_{j}^{q_{j,k}}i_{e}^{*}(v_{j,k})$
so $q_{j,k}=0$ and $i_{e}^{*}(v_{j,k})=T_{j,k}$ hence $v_{j,k}=Y_{j,k}$.
For similar reasons, $u_{j,k}=X_{j,k}$. We have proved that $\mu^{*}(T_{j,k})=X_{j,k}Y_{j,k}$ as desired.
9.3 A triangular representation
For $x,y\in V$, recall that we denote $\Delta_{x,y}:V\to V;v\mapsto x\circledast v\circledast y$. We consider the dual automorphism $\Delta^{*}_{x,y}\in{\mathrm{Aut}}_{\Gamma}(L_{V})$. The map $(x,y)\in V\times V\mapsto\Delta^{*}_{\check{x},y}$ is a representation of $V\times V$ in ${\mathrm{Aut}}_{\Gamma}(L_{V})$ (noting that we have used the inverse $\check{x}$ as the left multiplier). The following result says that, up to conjugacy, this representation has triangular image.
Theorem 9.4.
Let $\Gamma$ be a torsion-free hyperbolic group. Let $V$ be a connected algebraic group over $\Gamma$ with function group $L_{V}=L_{W}*{\mathbb{F}}_{r}$ where $L_{W}$ is the Grushko factor of $L_{V}$ containing $\Gamma$.
Then there exist some free factors $L_{V}^{r}=L_{V}\supset L_{V}^{r-1}\supset\dots\supset L_{V}^{0}=L_{W}$ and a finite index subgroup $V_{0}\subset V$ such that for all $i\leq r$, $L_{V}^{i}\simeq L_{V}^{i-1}*{\mathbb{Z}}$ and $L_{V}^{i}$ is invariant under $\Delta_{x,y}^{*}$ for all $x,y\in V_{0}$.
We will rely on the following result.
Theorem 9.5 ([HM20] [GH22, Theorem 1]).
Let $H\subset{\mathrm{Out}}(G_{1}*\dots*G_{k}*{\mathbb{F}}_{r})$ be a subgroup that preserves the conjugacy class of each $G_{i}$. Let ${\mathcal{F}}$ be a maximal $H$-periodic free factor system relative to $\{G_{1},\dots,G_{k}\}$.
If ${\mathcal{F}}$ is non-sporadic, then $H$ contains an element $\varphi$ preserving ${\mathcal{F}}$ and which is fully irreducible relative to ${\mathcal{F}}$.
If $G$ is a free product $G=G_{1}*\ldots*G_{k}*{\mathbb{F}}_{r}$, a free factor system for $G$ relative to $\{G_{1},\ldots,G_{k}\}$ is a collection =${\mathcal{F}}=\{[U_{1}],\ldots,[U_{p}]\}$ of conjugacy classes of non trivial proper subgroups $U_{1},\ldots,U_{p}$ of $G$ for which $G=U_{1}*\ldots*U_{p}*{\mathbb{F}}_{l}$, and such that each $G_{i}$ is contained in a conjugate of some $U_{j}$. Such a system is said to be sporadic if it corresponds to a decomposition of the form $G=U_{1}*U_{2}$ or $G=U_{1}*{\mathbb{F}}_{1}=U_{1}*{\mathbb{Z}}$.
We compare free factor systems as follows: $\{[U_{1}],\ldots,[U_{p}]\}\preceq\{[U^{\prime}_{1}],\ldots,[U^{\prime}_{p^{\prime}}]\}$ if for all $i\leq p$, there exists $j\leq p^{\prime}$ such that $U_{i}$ is conjugate into $U^{\prime}_{j}$. For example, a free factor system ${\mathcal{F}}$ of $G$ is relative to $\{[G_{1}],\ldots,[G_{k}]\}$ if $\{[G_{1}],\ldots,[G_{k}]\}\preceq{\mathcal{F}}$.
Proof of Theorem 9.4.
The natural map ${\mathrm{Aut}}_{\Gamma}(L_{V})\rightarrow{\mathrm{Out}}(L_{V})$ is injective (no nontrivial inner automorphism restricts to the identity on $\Gamma$), so we view ${\mathrm{Aut}}_{\Gamma}(L_{V})$ as a subgroup of ${\mathrm{Out}}(L_{V})$.
Consider the subgroup $H=\{\Delta_{x,y}^{*}\mid x,y\in V\}$ of ${\mathrm{Aut}}_{\Gamma}(L_{V})$. The group $H$ preserves $L_{W}$, so Theorem 9.5 applies to $H\subset{\mathrm{Out}}(L_{W}*{\mathbb{F}}_{r})$. Let ${\mathcal{F}}$ be a maximal $H$-periodic free factor system of $L_{V}$ relative to $L_{W}$. We denote by $H_{0}$ the finite index subgroup of $H$ which preserves ${\mathcal{F}}$.
We claim that ${\mathcal{F}}$ is sporadic. Otherwise, by Theorem 9.5, there exists $\varphi=\Delta^{*}_{u,v}\in H_{0}$ which is fully irreducible relative to ${\mathcal{F}}$.
Consider a train-track representative $f:T\rightarrow T$ of $\varphi$ where $T$ is a tree corresponding to some free splitting of $L_{V}$ relative to ${\mathcal{F}}$ [FM15], $\lambda>1$ its scaling factor, and $g\in L_{V}$ a $-{\mathcal{F}}$-peripheral element whose axis is $f$-legal. Then for all $n\in{\mathbb{N}}$, one has $||\varphi^{n}(g)||_{T}=\lambda^{n}||g||_{T}$. On the other hand, one has $\varphi^{n}(g)=\Delta_{u^{\circledast n},v^{\circledast n}}^{*}(g)$. Applying Corollary 7.8 to the space $Y=T$, we get that $||\Delta_{u^{\circledast n},v^{\circledast n}}^{*}(g)||_{T}$ is bounded by a constant $C_{g}$ which does not depend of $n$. This contradiction shows that ${\mathcal{F}}$ is sporadic, so that either $L_{V}=A*B$ with =${\mathcal{F}}=\{[A],[B]\}$ or $L_{V}=A*{\mathbb{Z}}$ with =${\mathcal{F}}=\{[A]\}$ with $\Gamma\subset A$ in both cases.
We now show that the former cannot hold. Consider a morphism $h\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ which kills $B$. The set $\{h\circ\alpha|\alpha\in H\}$ is discriminating by Lemma 4.4 since $\Delta^{*}_{1,v}=\rho^{*}_{v}$ lies in $H$ for all $v\in V$. By Lemma 5.5, the family $\{h\circ\alpha|\alpha\in H_{0}\}$ is still discriminating. Since all these morphisms kill $B$, this is a contradiction.
Therefore, one can write $L_{V}=L_{V}^{r-1}*{\mathbb{Z}}$ with =${\mathcal{F}}=\{[L_{V}^{r-1}]\}$ and $L_{W}\subset L_{V}^{r-1}$. We let $V_{r-1}$ be a finite index subgroup of $V$ such that $\Delta^{*}_{x,y}$ preserves $L_{V}^{r-1}$ up to conjugacy for all $x,y\in V_{r-1}$. The subgroup $V_{r-1}$ can be built as follows: consider the group morphism $\Delta:V\times V\to H$ defined by $(x,y)\mapsto\Delta^{*}_{u^{-1},v}$; then $\Delta^{-1}(H_{0})\subset V\times V$ projects to finite index subgroups of each of the two factors: take $V_{r-1}$ to be their intersection.
Since $\Delta^{*}_{x,y}$ is the identity on $\Gamma$, we note that $L_{V}^{r-1}$ is actually invariant (not only up to conjugacy).
We now iterate the argument, some care is needed because it is not a priori clear that the limit group $L_{V}^{r-1}$ corresponds to an algebraic group. Assume that we have constructed $L_{V}=L_{V}^{r}\supset L_{V}^{r-1}\supset\dots\supset L_{V}^{i}$ where for all $j\in\{i+1,\dots,r\}$, $L_{V}^{j}=L_{V}^{j-1}*{\mathbb{Z}}$, and for all $j\in\{i,\dots,r\}$, $L_{V}^{j}\supset L_{W}$ is invariant under $\Delta^{*}_{x,y}$ for all $x,y$ in a finite index subgroup $V_{j}\subset V$. Assume $L_{V}^{i}\neq L_{W}$.
Consider $H^{\prime}=\{(\Delta^{*}_{x,y})_{|L_{V}^{i}}\mid x,y\in V_{i}\}$. Let ${\mathcal{F}}^{\prime}$ be a maximal $H^{\prime}$-periodic free factor system of $L_{V}^{i}$ relative to $L_{W}$. As above, we claim that ${\mathcal{F}}^{\prime}$ is sporadic. If not, $H^{\prime}$ contains an automorphism $\varphi\in{\mathrm{Aut}}_{\Gamma}(L_{V}^{i})$ which is fully irreducible relative to ${\mathcal{F}}^{\prime}$. From $T^{\prime}$ a $L_{V}^{i}$-tree carrying a train track representative of $\varphi$, one can find a legal element $g\in A_{i}$ with $||\varphi^{n}(g)||_{T^{\prime}}\rightarrow\infty$ as $n$ goes to infinity.
One can extend the action of $L_{V}^{i}$ on $T^{\prime}$ to an action of $L_{V}$ on a larger tree $Y$: consider the graph of groups $T^{\prime}/L_{V}^{i}$ and add $r-i$ new loops with trivial edge group attached to the vertex whose vertex group contains $L_{W}$; one can take for $Y$ the Bass-Serre tree of the obtained graph of groups. By Corollary 7.8, $||\varphi^{n}(g)||_{Y}$ is bounded. Since $||\varphi^{n}(g)||_{Y}=||\varphi^{n}(g)||_{T^{\prime}}$, we get a contradiction showing that ${\mathcal{F}}^{\prime}$ is sporadic.
Arguing as above, ${\mathcal{F}}^{\prime}$ has to consist in a single factor =${\mathcal{F}}^{\prime}=\{[L_{V}^{i-1}]\}$ with $L_{V}^{i}=L_{V}^{i-1}*{\mathbb{Z}}$ and $L_{V}^{i-1}\supset L_{W}$ is invariant by $\Delta_{x,y}^{*}$ for $x,y$ in some finite index subgroup $V_{i-1}$ of $V$.
Iterating the argument at most $r$ times, we get the result. ∎
By Theorem 9.4, one can write $L_{V}$ as a free product $L_{V}=L_{W}*\langle T_{1},\dots,T_{r}\rangle$ where $\langle T_{1},\dots,T_{r}\rangle$ is free of rank $r$, and such that for all $i\leq r$, the subgroup $L_{V}^{i}=L_{W}*\langle T_{1},\dots,T_{i}\rangle$ is invariant under $\Delta^{*}_{x,y}$ for all $x,y$ in the finite index subgroup $V_{0}\subset V$.
Theorem 9.4 can be reformulated as follows:
Corollary 9.6.
Let $T_{1},\dots,T_{r}\in L_{V}$ be as above.
Then for all $i\leq r$, $x,y\in V_{0}$, the map $\Delta_{x,y}^{*}:L_{V}\rightarrow L_{V}$ acts as follows
$\Delta_{x,y}^{*}:T_{i}\ \mapsto\ g_{i,x,y}T_{i}^{\pm 1}g^{\prime}_{i,x,y}$
where $g_{i,x,y},g^{\prime}_{i,x,y}\in L_{V}^{i-1}=\langle L_{W},T_{1},\dots,T_{i-1}\rangle$.
9.4 Formal triangularization
We now interpret Corollary 9.6 in terms of the map $M^{*}:L_{V}\rightarrow L_{V^{3}}$ dual to $M:V^{3}\rightarrow V$ defined by $(x,z,y)\mapsto x\circledast z\circledast y$.
We denote by $T_{i},T_{j,k}\in L_{V}$ the coordinate system given by Proposition 9.3 and Corollary 9.6. We denote by $X_{i}$, $X_{j,k}$, $Y_{i}$, $Y_{j,k}$ (resp. $X_{i}$, $X_{j,k}$, $Z_{i}$, $Z_{j,k}$, $Y_{i}$, $Y_{j,k}$) the corresponding coordinate system on $L_{V^{2}}$ (resp. on $L_{V^{3}}$).
We denote by $L_{V}^{(X)}=\langle\Gamma,X_{jk},X_{i}|\;j\leq l,k\leq n_{j},i\leq r\rangle$ the copy of $L_{V}$ in $L_{V^{2}}$ or $L_{V^{3}}$ (according to the context) generated by the $X$ variables, and we define $L_{V}^{(Z)}$ and $L_{V}^{(Y)}$ similarly (where the $Z$ version only makes sense in $L_{V^{3}}$).
We then define the corresponding subgroups $L_{W}^{(X)}=\langle\Gamma,X_{jk}|\;j\leq l,k\leq n_{j}\rangle\subset L_{V}^{(X)}$, and $L_{V}^{(i,X)}=\langle L_{W}^{(X)},X_{1},\dots,X_{i}\rangle\subset L_{V}^{(X)}$ for $i\leq r$ the copies of $L_{W}$ and $L_{V}^{i}$ in $L_{V}^{(X)}$. We define $L_{W}^{(Z)}$, $L_{V}^{(i,Z)}$, $L_{W}^{(Y)}$, and $L_{V}^{(i,Y)}$ similarly. Thus, the subgroup $\langle L_{V}^{(X)},L_{V}^{(i-1,Z)},L^{(Y)}_{V}\rangle\subset L_{V^{3}}$ appearing in the next statement is the set of elements of $L_{V^{3}}$ that do not involve the generators $Z_{i},\dots,Z_{r}$.
Proposition 9.7.
The map $M^{*}:L_{V}\rightarrow L_{V^{3}}$ sends $T_{i}$ to $g_{i}Z_{i}{}^{\pm 1}g^{\prime}_{i}$ where the elements $g_{i},g^{\prime}_{i}\in L_{V^{3}}$ lie in $\langle L_{V}^{(X)},L_{V}^{(i-1,Z)},L^{(Y)}_{V}\rangle\subset L_{V^{3}}$.
One can think of this result as a formal version of Corollary 9.6, where the elements $g_{i,x,y}$ and $g^{\prime}_{i,x,y}$ (depending on $x$ and $y$) are now represented by fixed words on $X$ and $Y$. Additionally, the restriction $x,y\in V_{0}$ has vanished.
Given $x,y\in V$ we denote $i_{x,y}:V\to V^{3}$ the map defined by $v\mapsto(x,v,y)$. Its dual map $i^{*}_{x,y}:L_{V^{3}}\to L_{V}$ can be described as follows: if $h_{x},h_{y}\in\mathrm{Hom}_{\Gamma}(L_{V},\Gamma)$ denote the $\Gamma$-homomorphisms corresponding to $x$ and $y$, the restrictions of $i^{*}_{x,y}$ on $L_{V}^{(X)}$ and $L_{V}^{(Y)}$ coincides with $h_{x}$ and $h_{y}$ (with the natural identification) and its restriction to $L^{(Z)}_{V}$ is the identity (sending $Z_{i}$ to $T_{i}$ and $Z_{j,k}$ to $T_{j,k}$). Note for future use that this implies that $i^{*}_{x,y}$ maps the group $U=\langle L^{(X)}_{V},L^{(Z)}_{W},L^{(Y)}_{V}\rangle$ in $L_{W}$.
To prove Proposition 9.7 we will need the following lemma. Recall that $V_{0}$ is the finite index subgroup of $V$ from Theorem 9.4: for any $x,y\in V_{0}$, the map $\Delta_{x,y}^{*}$ is triangular.
Lemma 9.8.
The collection of $\Gamma$-homomorphisms $i^{*}_{x,y}:L_{V^{3}}\to L_{V}$ with $(x,y)\in V_{0}^{2}$ is a discriminating family.
Proof.
We view $L_{V^{3}}$ and $L_{V}$ as groups of functions on $V^{3}$ and $V$ with values in $\Gamma$.
Given $f\in L_{V^{3}}$ we have that $i^{*}_{x,y}(f)$ is the map $f(x,\cdot,y)$ so its image by $h_{v}$ is $f(x,v,y)$. This shows that $h_{v}\circ i^{*}_{x,y}=h_{(x,v,y)}$.
Since $V_{0}\times V\times V_{0}$ is a finite index subgroup of the connected algebraic group $V^{3}$, this is a dense subset of $V^{3}$ (Lemma 4.5) and by Corollary 2.12, the family of $\Gamma$-homomorphisms
$\{h_{(x,v,y)}|\,(x,v,y)\in V_{0}\times V\times V_{0}\}$
is discriminating. Since $h_{(x,v,y)}$ factors through $i^{*}_{x,y}$ this concludes the proof. ∎
Proof of Proposition 9.7.
Denote $U=\langle L^{(X)}_{V},L^{(Z)}_{W},L^{(Y)}_{V}\rangle\subset L_{V^{3}}$. Write $L_{V^{3}}$ as a free product $U*{\mathbb{F}}_{r}$ where ${\mathbb{F}}_{r}$ is the free group ${\mathbb{F}}_{r}=\langle Z_{1},\dots,Z_{r}\rangle$. Write $M^{*}(T_{i})$ as a reduced word in this free product: $M^{*}(T_{i})=u_{0}w_{1}u_{1}w_{2}\dots u_{n}w_{n+1}$ with $u_{j}\in U,w_{j}\in{\mathbb{F}}_{r}$ and $u_{j},w_{j}$ non-trivial except maybe for $u_{0},w_{n+1}$. It suffices to prove that $w_{j}\in\langle Z_{1},\dots,Z_{i}\rangle$ for all $j$, and that $Z_{i}^{\pm 1}$ appears exactly once in exactly one of the $w_{j}$’s.
By Lemma 9.8, there exist $x,y\in V_{0}^{2}$ such that for all $j$, the element $u^{\prime}_{j}=i^{*}_{(x,y)}(u_{j})\in L_{V}$ is non-trivial whenever $u_{j}$ is non-trivial. As noticed above, $i_{x,y}^{*}(U)\subset L_{W}$ so $u^{\prime}_{j}\in L_{W}$. Now
$i^{*}_{(x,y)}\circ M^{*}(T_{i})=u^{\prime}_{0}w_{1}u^{\prime}_{2}w_{2}\dots u^{\prime}_{n}w_{n+1}\in L_{W}*\langle T_{1},\dots,T_{r}\rangle$
(under the natural identification of $\langle Z_{1},\dots,Z_{r}\rangle\subset L^{(Z)}_{V}$ with its copy $\langle T_{1},\dots,T_{r}\rangle\subset L_{V}$). This is a reduced word in the free product $L_{V}=L_{W}*\langle T_{1},\dots,T_{r}\rangle$.
On the other hand, we have $\Delta_{x,y}=M\circ i_{(x,y)}$ so $\Delta^{*}_{x,y}=i^{*}_{(x,y)}\circ M^{*}$ so by Corollary 9.6,
$i^{*}_{(x,y)}\circ M^{*}(T_{i})=g_{i,x,y}T_{i}^{\pm 1}g^{\prime}_{i,x,y}$
where $g_{i,x,y},g^{\prime}_{i,x,y}\in L_{V}^{i-1}=L_{W}*\langle T_{1},\dots,T_{i-1}\rangle$. This shows that $w_{1},\ldots,w_{n+1}$ lie in $\langle T_{1},\ldots,T_{i}\rangle$ (i. e. the generators $T_{i+1},\dots T_{r}$ do not appear) and that $T_{i}^{\pm 1}$ appears exactly once in exactly one of the $w_{j}$.
This proves the proposition. ∎
9.5 Structure of the multiplication law
Recall that $\mu:V\times V\rightarrow V$ is the multiplication law, and $\mu^{*}:L_{V}\rightarrow L_{V^{2}}$ is the dual map.
We have already seen that the multiplication is standard in restriction to the abelian generators of $L_{W}\subset L_{V}$ (see Proposition 9.3).
We now turn to the description of $\mu^{*}$ on the free generators $T_{1},\ldots,T_{r}$ of $L_{V}$. For $i\leq r$, we denote by $L_{V^{3}}^{i}=\langle L_{W^{3}},X_{1},Z_{1},Y_{1},\dots,X_{i},Z_{i},Y_{i}\rangle$ and $L_{V^{2}}^{i}=\langle L_{W^{2}},X_{1},Y_{1}\dots,X_{i},Y_{i}\rangle$, i. e. an element lies in $L_{V^{3}}^{i}$ (resp. $L_{V^{2}}^{i}$) if it does not involve the free generators with index greater than $i$.
Proposition 9.9.
For all $i\leq r$, there exists $a_{i},b_{i},c_{i}\in L_{V^{2}}^{i-1}$ such that either
$\mu^{*}(T_{i})=a_{i}\cdot\left(X_{i}\right)^{\pm 1}\cdot b_{i}\cdot\left(Y_{i}\right)^{\pm 1}\cdot c_{i}$
or
$\mu^{*}(T_{i})=a_{i}\cdot\left(Y_{i}\right)^{\pm 1}\cdot b_{i}\cdot\left(X_{i}\right)^{\pm 1}\cdot c_{i}.$
Proof.
Since $\mu(x,y)=M(x,y,e)=M(e,x,y)$, one has $\mu^{*}=i_{e}^{*}\circ M^{*}=i_{e}^{\prime*}\circ M^{*}$ where $i_{e},i^{\prime}_{e}:V^{2}\rightarrow V^{3}$ are defined by $i_{e}(x,y)=(x,y,e)$ and $i_{e}^{\prime}(x,y)=(e,x,y)$. The dual morphisms $i_{e}^{*}$ and $i_{e}^{\prime*}$ satisfy
$i_{e}^{*}:L_{V^{3}}$ $\rightarrow L_{V^{2}}$ $i_{e}^{\prime*}:L_{V^{3}}$ $\rightarrow L_{V^{2}}$
$X_{i},X_{j,k}$ $\mapsto X_{i},X_{j,k}$ $X_{i},X_{j,k}$ $\mapsto h_{e}(X_{i}),h_{e}(X_{j,k})=1\in\Gamma$
$Z_{i},Z_{j,k}$ $\mapsto Y_{i},Y_{j,k}$ $Z_{i},Z_{j,k}$ $\mapsto X_{i},X_{j,k}$
$Y_{i},Y_{j,k}$ $\mapsto h_{e}(Y_{i}),h_{e}(Y_{j,k})=1\in\Gamma$ $Y_{i},Y_{j,k}$ $\mapsto Y_{i},Y_{j,k}$
Fix $i\leq r$ and recall that $L_{V^{2}}=L_{V^{2}}^{i-1}*\langle X_{i},Y_{i},\dots,X_{r},Y_{r}\rangle$. Write $\mu^{*}(T_{i})$ as a reduced word as follows:
$\mu^{*}(T_{i})=g_{0}S_{1}g_{1}\dots S_{n}g_{n}$ (1)
where for each $j\leq n$, $S_{j}$ is a letter in $\{X_{i}^{\pm 1},Y_{i}^{\pm 1},\dots,X_{r}^{\pm 1},Y_{r}^{\pm 1}\}$, and $g_{j}\in L_{V^{2}}^{i-1}$, and $g_{j}$ is nontrivial when $S_{j+1}=S_{j}^{-1}$.
By Proposition 9.7 we have
$M^{*}(T_{i})=g_{i}{Z_{i}}^{\pm 1}g^{\prime}_{i}\text{ for some }g_{i},g^{\prime}_{i}\in\langle L_{V^{3}}^{i-1},X_{i},Y_{i}\dots,X_{r},Y_{r}\rangle.$
Thus
$\mu^{*}(T_{i})=i^{*}_{e}\circ M^{*}(T_{i})=i_{e}^{*}(g_{i}){Y_{i}}^{\pm 1}i_{e}^{*}(g^{\prime}_{i})$
with $i_{e}^{*}(g_{i}),i_{e}^{*}(g^{\prime}_{i})\in\langle L_{V^{2}}^{i-1},X_{i},\dots,X_{r}\rangle$. This shows that the letters $Y_{i+1}^{\pm 1},\dots,Y_{r}^{\pm 1}$ don’t occur in (1), and that $Y_{i}^{\pm 1}$ occurs exactly once.
By a similar argument using $i_{e}^{\prime*}$, we see that the letters $X_{i+1}^{\pm 1},\dots,X_{r}^{\pm 1}$ don’t occur in (1), and that $X_{i}^{\pm 1}$ occurs exactly once. This concludes the proof. ∎
The following lemma says that we may assume that the exponents of $X_{i},Y_{i}$ in Proposition 9.9 are positive and that $X_{i}$ comes before $Y_{i}$.
Lemma 9.10.
Up to changing $T_{i}$ to $T_{i}^{-1}$ and $X_{i},Y_{i}$ to $X_{i}^{-1},Y_{i}^{-1}$ accordingly (see Remark 9.1), there exists $a_{i},b_{i},c_{i}\in L_{V^{2}}^{i-1}$ such that
$\mu^{*}(T_{i})=a_{i}X_{i}b_{i}Y_{i}c_{i}.$
Proof.
We apply Proposition 9.9. If $\mu^{*}(T_{i})=a_{i}Y_{i}^{\varepsilon}b_{i}X_{i}^{\varepsilon^{\prime}}c_{i}$, for some $\varepsilon,\varepsilon^{\prime}\in\{\pm 1\}$ and $a_{i},b_{i},c_{i}\in L_{V^{2}}^{i-1}$, then we change $T_{i}$ to $T^{\prime}_{i}=T_{i}^{-1}$ and $X_{i},Y_{i}$ to $X^{\prime}_{i}=X_{i}^{-1},Y^{\prime}_{i}=Y_{i}^{-1}$ accordingly. This reverses the order of appearance of $X_{i}$ and $Y_{i}$: we get $\mu^{*}(T^{\prime}_{i})=c_{i}^{-1}X_{i}^{\prime}{}^{\varepsilon^{\prime}}b_{i}^{-1}Y_{i}^{\prime\varepsilon}a_{i}^{-1}$. Thus, we may assume that
$\mu^{*}(T_{i})=a_{i}X_{i}^{\varepsilon}b_{i}Y_{i}^{\varepsilon^{\prime}}c_{i}.$
We use the fact that $\mu(v,e)=v$ for all $v\in V$, i. e. $\mu\circ i_{e}=\mathrm{id}_{V}$ where $i_{e}:V\rightarrow V^{2}$ is defined by $i_{e}(v)=(v,e)$. This translates into $i_{e}^{*}\circ\mu^{*}=\mathrm{id}_{L_{V}}$. Similarly, we have $i_{e}^{\prime*}\circ\mu^{*}=\mathrm{id}_{L_{V}}$ where $i^{\prime}_{e}(v)=(e,v)$. Note that
$i_{e}^{*}:L_{V^{2}}$ $\rightarrow L_{V}$ $i_{e}^{\prime*}:L_{V^{2}}$ $\rightarrow L_{V}$
$X_{i},X_{j,k}$ $\mapsto T_{i},T_{j,k}$ $X_{i},X_{j,k}$ $\mapsto h_{e}(X_{i}),h_{e}(X_{j,k})=1\in\Gamma$
$Y_{i},Y_{j,k}$ $\mapsto h_{e}(Y_{i}),h_{e}(Y_{j,k})=1\in\Gamma$ $Y_{i},Y_{j,k}$ $\mapsto T_{i},T_{j,k}$
We have that $i_{e}^{*}\circ\mu^{*}(T_{i})=i_{e}^{*}(a_{i})\cdot T_{i}^{\varepsilon}\cdot i_{e}^{*}(b_{i}Y_{i}^{\varepsilon^{\prime}}c_{i})$. Since this must be equal to $T_{i}$ and since $i_{e}^{*}(a_{i}),i_{e}^{*}(b_{i}Y_{i}^{\varepsilon^{\prime}}c_{i})\in L_{V}^{i-1}$, the exponent $\varepsilon$ has to be positive. Arguing with $i_{e}^{\prime*}$ instead of $i_{e}^{*}$, we similarly get $\varepsilon^{\prime}=+1$ which concludes the proof. ∎
From now on, we take $a_{i},b_{i},c_{i}\in L_{V^{2}}^{i-1}$ as in Lemma 9.10. The following proposition is the last step in the proof of the main result. In this statement, given $u\in L_{V}$, we denote by $u^{(X)}\in L_{V}^{(X)}\subset L_{V^{2}}$, $u^{(Y)}\in L_{V}^{(Y)}\subset L_{V^{2}}$ the images of $u$ under the embeddings of $L_{V}$ into $L_{V^{2}}$.
Proposition 9.11.
For each $i$ there exist $u_{i},v_{i}\in L_{V}^{i-1}$ such that if we define $T^{\prime}_{i}=u_{i}T_{i}v_{i}$, and change accordingly $X_{i},Y_{i}$ to $X^{\prime}_{i}=u_{i}^{(X)}X_{i}v_{i}^{(X)}$ and $Y^{\prime}_{i}=u_{i}^{(Y)}Y_{i}v_{i}^{(Y)}$ (see Remark 9.1), we have
$\mu^{*}(T^{\prime}_{i})=X_{i}^{\prime}Y_{i}^{\prime}.$
We will need two lemmas before proving the proposition.
Given $u\in L_{V^{2}}$, we denote by $u^{(XZ)}\in\langle L_{V}^{(X)},L_{V}^{(Z)}\rangle\subset L_{V^{3}}$ the image of $u$ under the embeddings of $L_{V^{2}}$ into $L_{V^{3}}$ sending respectively $X_{i},X_{jk},Y_{i},Y_{jk}$ to $X_{i},X_{jk},Z_{i},Z_{jk}$. Similarly we denote by $u^{(ZY)}\in\langle L_{V^{3}}^{(Z)},L_{V^{3}}^{(Y)}\rangle$ its image under the embedding sending respectively $X_{i},X_{jk},Y_{i},Y_{jk}$ to $Z_{i},Z_{jk},Y_{i},Y_{jk}$.
Define $\theta_{L},\theta_{R}:V^{3}\rightarrow V^{2}$ by $\theta_{L}(x,y,z)=(x\circledast y,z)$ and $\theta_{R}(x,y,z)=(x,y\circledast z)$. The dual maps satisfy
$\theta_{L}^{*}:L_{V^{2}}$ $\rightarrow L_{V^{3}}$ $\theta_{R}^{*}:L_{V^{2}}$ $\rightarrow L_{V^{3}}$
$X_{i}$ $\mapsto a_{i}^{(XZ)}X_{i}b_{i}^{(XZ)}Z_{i}c_{i}^{(XZ)}$ $X_{i}$ $\mapsto X_{i}$
$Y_{i}$ $\mapsto Y_{i}$ $Y_{i}$ $\mapsto a_{i}^{(ZY)}Z_{i}b_{i}^{(ZY)}Y_{i}c_{i}^{(ZY)}.$
In particular, we have that $\theta_{L}^{*}(L_{V^{2}}^{i-1})\subset L_{V^{3}}^{i-1}$, and $\theta_{R}^{*}(L_{V^{2}}^{i-1})\subset L_{V^{3}}^{i-1}$.
Associativity says that $\mu\circ\theta_{L}=\mu\circ\theta_{R}$, i. e. $\theta_{L}^{*}\circ\mu^{*}=\theta_{R}^{*}\circ\mu^{*}$. This is the basis of the following relations.
Lemma 9.12.
With $a_{i},b_{i},c_{i}$ as in Lemma 9.10 we have in $L_{V^{3}}$
$\theta_{R}^{*}(a_{i})$ $=\theta_{L}^{*}(a_{i})a^{(XZ)}_{i}$ (2)
$b^{(XZ)}_{i}$ $=\theta_{R}^{*}(b_{i})a^{(ZY)}_{i}$ (3)
$b^{(ZY)}_{i}$ $=c^{(XZ)}_{i}\theta_{L}^{*}(b_{i})$ (4)
$\theta_{L}^{*}(c_{i})$ $=c^{(ZY)}_{i}\theta_{R}^{*}(c_{i})$ (5)
Proof.
Recall that we have $\mu^{*}(T_{i})=a_{i}X_{i}b_{i}Y_{i}c_{i}\in L_{V^{2}}$, where $a_{i},b_{i},c_{i}\in L_{V^{2}}^{i-1}$ (i. e. do not involve $X_{i},Y_{i},\dots,X_{r},Y_{r}$). Thus
$\theta_{L}^{*}(\mu^{*}(T_{i}))=\theta_{L}^{*}(a_{i})\cdot a_{i}^{(XZ)}X_{i}b_{i}^{(XZ)}Z_{i}c_{i}^{(XZ)}\cdot\theta_{L}^{*}(b_{i})\cdot Y_{i}\cdot\theta_{L}^{*}(c_{i})$
and
$\theta_{R}^{*}(\mu^{*}(T_{i}))=\theta_{R}^{*}(a_{i})\cdot X_{i}\cdot\theta_{R}^{*}(b_{i})\cdot a_{i}^{(ZY)}Z_{i}b_{i}^{(ZY)}Y_{i}c_{i}^{(ZY)}\cdot\theta_{R}^{*}(c_{i}).$
We want to identify the two expressions using the free product structure
$L_{V^{3}}^{i-1}*\langle X_{i},Z_{i},Y_{i},\ldots,X_{r},Z_{r},Y_{r}\rangle.$
Since $a_{i},b_{i},c_{i}\in L_{V^{2}}^{i-1}$, their copies in $L_{V^{3}}^{(XZ)}$ and $L_{V^{3}}^{(ZY)}$ are in $L_{V^{3}}^{i-1}$ i. e. don’t involve the free generators with index at least $i$. Since $\theta_{L}^{*}(L_{V^{2}}^{i-1})\subset L_{V^{3}}^{i-1}$ and $\theta_{R}^{*}(L_{V^{2}}^{i-1})\subset L_{V^{3}}^{i-1}$, the same holds for the images of $a_{i},b_{i},c_{i}$ by $\theta_{L}^{*},\theta_{R}^{*}$. By identification, we get the desired four equations. ∎
Lemma 9.13.
Fix $i_{0}\leq r$. Assume that $\mu^{*}(T_{i})=X_{i}Y_{i}$ for all $i<i_{0}$.
With the notations of Lemma 9.10, we have $b_{i_{0}}\in(L_{V}^{({i_{0}}-1,X)}).(L_{V}^{({i_{0}}-1,Y)})$.
Before proving the lemma, recall that we write $L_{V}$ as the fundamental group of a graph of groups $\Lambda$ as in Theorem 8.1. Recall that the abelian vertex groups of $\Lambda$ are denoted by $A_{j}$, and that we write $A_{j}=\langle c_{j}\rangle\oplus N_{j}$.
Following [Ser77], we define a reduced word over $\Lambda$ to be a word of the form
$\gamma_{0}x_{1}\gamma_{1}\ldots\gamma_{n-1}x_{n}\gamma_{n}$
with $\gamma_{k}\in\Gamma$ and $x_{k}\in\{T_{1}^{\pm 1},\ldots,T_{r}^{\pm 1}\}\sqcup\cup_{j}A_{j}$ satisfying the following conditions:
  • if $x_{k}\in A_{j}$, then $x_{k}\notin\langle c_{j}\rangle$;
  • if $x_{k},x_{k+1}\in A_{j}$ then $\gamma_{k}\notin\langle c_{j}\rangle$;
  • if $x_{k}=T_{i}^{\varepsilon}$ and $x_{k+1}=T_{i}^{-\varepsilon}$ then $\gamma_{k}\neq 1$.
If two reduced words $\gamma_{0}x_{1}\gamma_{1}\ldots\gamma_{n-1}x_{n}\gamma_{n}$ and $\gamma^{\prime}_{0}x^{\prime}_{1}\gamma^{\prime}_{1}\ldots\gamma^{\prime}_{n^{\prime}-1}x^{\prime}_{n^{\prime}}\gamma^{\prime}_{n^{\prime}}$ represent the same element, then by [Ser77, Exercice §5.2] the following hold: $(i)$ $n=n^{\prime}$ and $(ii)$ for each $k$, if $x_{k}=T_{i}^{\varepsilon}$, then $x^{\prime}_{k}=x_{k}$, and if $x_{k}\in A_{j}$, then $x^{\prime}_{k}=x_{k}c_{j}^{q}$ for some $q\in{\mathbb{Z}}$.
Analogous statements hold for reduced elements in $L_{V^{2}}$ and $L_{V^{3}}$ with respect to the graphs of groups $\Lambda_{V^{2}}$, $\Lambda_{V^{3}}$.
Proof of Lemma 9.13.
Write $b_{i_{0}}\in L_{V^{2}}^{i_{0}-1}$ as a reduced word $b_{i_{0}}=\gamma_{0}x_{1}\gamma_{1}\ldots x_{n}\gamma_{n}$ over $\Lambda_{V^{2}}$: here, the $\gamma_{k}$ are in $\Gamma$ and the $x_{k}$ are in $\{X_{1}^{\pm 1},Y_{1}^{\pm 1},\ldots,X_{i_{0}-1}^{\pm 1},Y_{i_{0}-1}^{\pm 1}\}\sqcup\cup_{j}A^{(XY)}_{j}$.
It suffices to show that there is no $Y$’s before the $X$’s in this word. More formally, say that some element $u\in L_{V^{2}}$ does not involve the $X$’s (resp. the $Y$’s) if $u\in L_{V}^{(Y)}$ (resp. $u\in L_{V}^{(X)}$) (we will use similar wordings in $L_{V^{3}}$). We assume by contradiction that there are $k_{1}<k_{2}$ for which $x_{k_{1}}\notin L_{V}^{(X)}$ and $x_{k_{2}}\notin L_{V}^{(Y)}$ and we will show that this contradicts Equation 3 for $i=i_{0}$.
When $x_{k}$ lies in some abelian vertex group $A_{j_{k}}^{(XY)}=\langle c_{j_{k}}\rangle\oplus N_{j_{k}}^{(X)}\oplus N_{j_{k}}^{(Y)}$ (for some $j_{k}\leq l$), we decompose $x_{k}$ accordingly as
$x_{k}=c_{j_{k}}^{q_{k}}u_{k}u_{k}^{\prime}$
with $u_{k}\in N_{j_{k}}^{(X)}$, $u_{k^{\prime}}\in N_{j_{k}}^{(Y)}$ not both trivial. Then by Proposition 9.3,
$\theta_{R}^{*}(x_{k})=c_{j_{k}}^{q_{k}}u_{k}^{(X)}u_{k}^{\prime(Z)}u_{k}^{\prime(Y)}\in A_{j}^{V^{3}}$ (6)
which does not lie in $\langle c_{j_{k}}\rangle$.
We first claim that the decomposition $\theta_{R}^{*}(b_{i_{0}})=\gamma_{0}\theta_{R}^{*}(x_{1})\gamma_{1}\ldots\theta_{R}^{*}(x_{n})\gamma_{n}$ is a reduced word in $L_{V^{3}}$. It suffices to check that every subword of the form $w=\theta_{R}^{*}(x_{k})\gamma_{k}\theta_{R}^{*}(x_{k+1})$ is reduced.
If $x_{k}$ and $x_{k+1}$ lie in distinct abelian vertex groups, then $w$ is reduced by (6). If $x_{k}$ and $x_{k+1}$ lie in the same abelian vertex group $A^{(XY)}_{j_{k}}$, then $\gamma_{k}\notin\langle c_{j_{k}}\rangle$ and $w$ is reduced.
If $x_{k}$ and $x_{k+1}$ are free letters; then $\gamma_{k}\neq 1$ if $x_{k+1}=x_{k}^{-1}$. Since by assumption, for all $i<i_{0}$
$\theta_{R}^{*}(X_{i})=X_{i},\qquad\theta_{R}^{*}(Y_{i})=Z_{i}Y_{i}$ (7)
one easily deduces that $w$ is reduced.
Finally, if $x_{k}$ is a free letter and $x_{k+1}$ lies in some abelian vertex group or vice versa, then $w$ is clearly reduced, which proves our claim.
Since $x_{k_{1}}$ involves $Y$’s and $x_{k_{2}}$ involves $X$’s, the reduced word for $\theta_{R}^{*}(b_{0})$ has a letter involving $Y$’s before a letter involving $X$’s by (6) and (7).
When reducing $\theta_{R}^{*}(b_{0})a_{i}^{(ZY)}$, the reduction will not cancel the letters involving the $X$’s so the leftmost letter involving $Y$’s in $\theta_{R}^{*}(b_{0})$ will not cancel. Now (3), together with the results on reduced words representing the same element of [Ser77] referred to above, implies that $\theta_{R}^{*}(b_{0})a_{i}^{(ZY)}\in L_{V^{3}}^{(XZ)}$, a contradiction. ∎
Proof of Proposition 9.11.
By induction, we may assume that the result holds for the indices $11,\ldots,i-$ 1. Up to a change of variables, we may thus assume that $\mu^{*}(T_{1})=X_{1}Y_{1},\ldots,\mu^{*}(T_{i-1})=X_{i-1}Y_{i-1}$.
By Lemma 9.13, there exist $b^{\prime}_{i},b^{\prime\prime}_{i}\in L_{V}^{i-1}$, such that $b_{i}=b_{i}^{\prime(X)}.b_{i}^{\prime\prime(Y)}$ (where, as usual, notations such as $u^{(X)}$ denotes the copy of $u\in L_{V}$ in $L_{V}^{(X)}$).
Equation (3) gives
$b_{i}^{\prime(X)}b^{\prime\prime(Z)}_{i}=\theta_{R}^{*}\left(b_{i}^{\prime(X)}b_{i}^{\prime\prime(Y)}\right)a^{(ZY)}_{i}=b_{i}^{\prime(X)}\Big{[}\mu^{*}(b_{i}^{\prime\prime})\Big{]}^{(ZY)}a^{(ZY)}_{i}$
hence $b^{\prime\prime(Z)}_{i}=\Big{[}\mu^{*}(b_{i}^{\prime\prime})\Big{]}^{(ZY)}a^{(ZY)}_{i}$. Both sides of this equation live in the subgroup $L^{(ZY)}_{V^{2}}\subset L_{V^{3}}$ isomorphic to $L_{V^{2}}$, so one can rewrite this equation in $L_{V^{2}}$ as
$b_{i}^{\prime\prime(X)}=\mu^{*}(b^{\prime\prime}_{i})a_{i}$ (8)
Similarly, we using (4), we get
$b_{i}^{\prime(Y)}=c_{i}\mu^{*}(b^{\prime}_{i}).$ (9)
Define $T^{\prime}_{i}=b^{\prime\prime}_{i}T_{i}b^{\prime}_{i}$ in $L_{V}$, and $X^{\prime}_{i}=b_{i}^{\prime\prime(X)}X_{i}b_{i}^{\prime(X)}$, $Y^{\prime}_{i}=b_{i}^{\prime\prime(Y)}X_{i}b_{i}^{\prime(Y)}$ in $L_{V^{2}}$. We have
$\mu^{*}(T^{\prime}_{i})$ $=\mu^{*}(b^{\prime\prime}_{i})\cdot a_{i}X_{i}b_{i}^{\prime(X)}b_{i}^{\prime\prime(Y)}Y_{i}c_{i}\cdot\mu^{*}(b^{\prime}_{i})$
$=\mu^{*}(b^{\prime\prime}_{i})a_{i}[b_{i}^{\prime\prime(X)}]^{-1}X_{i}^{\prime}Y_{i}^{\prime}[b_{i}^{\prime(Y)}]^{-1}c_{i}\mu^{*}(b^{\prime}_{i})$
$=X_{i}^{\prime}Y_{i}^{\prime}$
using (8) and (9). ∎